### v1

# Submodular Maximization over a Matroid $k$-Intersection: Multiplicative Improvement over Greedy

Abstract
We study the problem of maximizing a non-negative monotone submodular objective $f$ subject to the intersection of $k$ arbitrary matroid constraints. The natural greedy algorithm guarantees $(k+1)$-approximation for this problem, and the state-of-the-art algorithm only improves this approximation ratio to $k$. We give a $\frac{2k\ln 2}{1+\ln 2}+O(\sqrt{k})<0.819k+O(\sqrt{k})$ approximation algorithm for this problem. Our result is the first multiplicative improvement over the approximation ratio of the greedy algorithm for general $k$. We further show that our algorithm can be used to obtain roughly the same approximation ratio also for the more general problem in which the objective is not guaranteed to be monotone (the sublinear term in the approximation ratio becomes $O(k^{2/3})$ rather than $O(\sqrt{k})$ in this case).
All of our results hold also when the $k$-matroid intersection constraint is replaced with a more general matroid $k$-parity constraint. Furthermore, unlike the case in many of the previous works, our algorithms run in time that is independent of $k$ and polynomial in the size of the ground set. Our algorithms are based on a hybrid greedy local search approach recently introduced by Singer and Thiery [38] for the weighted matroid $k$-intersection problem, which is a special case of the problem we consider. Leveraging their approach in the submodular setting requires several non-trivial insights and algorithmic modifications since the marginals of a submodular function $f$, which correspond to the weights in the weighted case, are not independent of the algorithm’s internal randomness. In the special weighted case studied by [38], our algorithms reduce to a variant of the algorithm of [38] with an improved approximation ratio of $k\ln 2+1-\ln 2+\varepsilon<0.694k+0.307+\varepsilon$, compared to an approximation ratio of $\frac{k+1}{2\ln 2}\approx 0.722k+0.722$ guaranteed by Singer and Thiery [38].
keywords: submodular function, matroid $k$-parity, matroid intersection, local search, greedy
1 Introduction
An important branch of combinatorial optimization studies the maximization of linear and submodular functions subject to natural families of combinatorial constraints. One large such family is matroid $k$-intersection—the family of constraints that can be represented as the intersection of $k$ matroid constraints. In $1978$, Fisher, Nemhauser and Wolsey [15] showed that the natural greedy algorithm guarantees $(k+1)$-approximation for the problem of maximizing a monotone submodular function subject to a matroid $k$-intersection constraint, and for linear functions the above approximation ratio improves to $k$ [25, 22].
Improving over the above guarantees of the greedy algorithm has been an open question for many years, and until recently success has mostly been limited to special cases such as $k$-dimensional matching constraints. Here, we are given a hypergraph $H=(V,E)$ such that the set $V$ of vertices can be partitioned into $k$ disjoint subsets (or parts) $V_{1},V_{2}\dotsc,V_{k}$, and each hyperedge $e\in E$ contains at most 11 vertex from each one of the subsets $V_{i}$. Such hypergraphs are known as $k$-partite hypergraphs. The $k$-dimensional matching constraint requires that we select a set of vertex disjoint hyperedges from $E$ (i.e., a matching in $H$). This constraint can be captured by an intersection of $k$ partition matroid constraints, each enforcing that no vertex of a particular set $V_{i}$ is contained in two hyperedges.
Following a long line of works (including [1, 7, 3, 4, 34, 36, 40]), Neuwohner [35] obtained an approximation ratio of $0.4986k+O(1)$ for weighted $k$-dimensional matching (i.e., the problem of maximizing a linear function subject to a $k$-dimensional matching constraint). [^1] We note that Thiery [41] derived better approximation results for small to moderate values of $k$, but for large $k$ values he only obtains a weaker guarantee of $0.5k$. In the more general case of monotone submodular objectives, Ward [42] obtained a slightly weaker approximation ratio of $k/2+O(1)$. In the less general case of a cardinality objective (i.e., when the goal is to maximize the number of elements in the solution), the best-known approximation is $k/3+O(1)$ [39, 11, 16]. On the hardness side, Lee, Svensson and Thiery [27] have shown that even in the unweighted case, no polynomial time algorithm can obtain better than $k/12$-approximation, improving on a previous lower bound of $\Omega(k/\log k)$ due to Hazan, Safra and Schwartz [19].
All the above algorithmic results apply even for the more general case of $k$-set packing constraints, which are similar to $k$-dimensional matching constraints, but allow the graph $H$ to be a general $k$-uniform hypergraph rather than a $k$-partite hypergraph. Unlike $k$-dimensional matching constraints, $k$-set packing constraints are not captured by matroid $k$-intersection, but they have a different crucial property, namely, that feasible individual exchanges can be carried out together as a single exchange. This property holds for partition matroids and, more generally, to strongly base orderable matroids, but does not necessarily hold even for a single general matroid constraint.
For general matroid $k$-intersection constraints, Lee, Sviridenko and Vondrák [28] gave a local search algorithm with an approximation ratio of k1k-1 for linear functions and $k$ for monotone submodular functions. In the special case of a cardinality objective, the same authors [29] gave a different local search algorithm with an approximation ratio of $k/2+\varepsilon$. We note that while the improvement in approximation for the cardinality objective has been by a multiplicative factor of 22, the improvements for submodular and linear functions have, until recently, been only by an additive term. This gap stems from the fact that the analysis presented in [29] is based on global arguments that are appropriate for the unweighted case in which all elements are equally valuable, but obtaining a result for a weighted problem requires more local arguments.
Recently, Singer and Thiery [38] have overcome this obstacle to give a $\frac{k}{2\ln 2}+O(1)\approx 0.722k+O(1)$-approximation for linear objective functions by combining greedy and local search techniques. Their algorithm randomly partitions elements into weight classes, and processes these classes in descending order of weight, similar to the greedy algorithm. In contrast to the standard greedy algorithm, however, the algorithm uses a local search procedure to find a good set of elements to add when processing each weight class. Intuitively, since the elements within each weight class have roughly equal weights, performing local search within classes allows the algorithm to make use of improved guarantees from the unweighted setting, while processing the classes greedily allows the algorithm to avoid the worst-case local optima of a simple global local search procedure.
As we have noted, matroid $k$-intersection and $k$-set packing constraints are two incomparable generalizations of $k$-dimensional matching constraints. They can be unified under the family of matroid $k$-parity constraints. Both the results of Lee, Sviridenko and Vondrák [29] for cardinality objectives and of Singer and Thiery [38] for linear objectives hold even for this general family of constraints. See Figure 1 for a graphical depiction of the relationships between the different families of constraints we have discussed. The figure also states the state-of-the-art result for each combination of constraint family and class of objective functions.
$k$-Dimensional Matchingkk-Set Packing
C: $k/3+O(1)$ [39] L: $0.4986k+O(1)$ [35]
MS: $k/2+O(1)$ [42] NS: $0.819k+O(k^{2/3})$
  NS: (prev. $k+O(1)$ [14])
$k$-Intersection
C: $k/2+\varepsilon$ [29] L: $0.694k+O(1)$
  L: (prev. $0.722k$ [38])
MS: $0.819k+O(\sqrt{k})$ NS: $0.819k+O(k^{2/3})$
  MS: (prev. $k+\varepsilon$ [28])   NS: (prev. $k+O(1)$ [28])
Matroid $k$-Parity
C: $k/2+\varepsilon$ [29] L: $0.694k+O(1)$
  L: (prev. $0.722k$ [38])
MS: $0.819k+O(\sqrt{k})$ NS: $0.819k+O(k^{2/3})$
  MS: (prev. $k+1$ [15, 9])   NS: (prev. $k+O(\sqrt{k})$ [13])
Figure 1: The four families of constraints discussed in Section 1. An arrow from family A of constraints to family B indicates that family B generalizes family A. Below each constraint family, we list the state-of-the-art approximation ratios for maximizing cardinality (C), linear (L), monotone submodular (MS) and/or non-monotone submodular (NS) objectives subject to constraints of this family. New results of this paper are marked with a grey background (and the results they improve over appear below them in parentheses). To avoid repetition, the approximation ratios for $k$-Dimensional Matching constraints are omitted as they are identical to the ratios for the more general $k$-Set Packing constraints.
1.1 Our Results
In this paper, we obtain the first result that multiplicatively improves over the approximation guarantee of the greedy algorithm for the problem of maximizing a monotone submodular objective function subject to a general matroid $k$-intersection constraint. Our technique is based on the one of Singer and Thiery [38], and like theirs, works also for the more general family of matroid $k$-parity constraints. However, the submodular case that we consider introduces several additional difficulties that must be overcome.
Previous works [28, 42, 6, 8] have adapted local search algorithms for weighted maximization to submodular objectives by using the marginal or incremental contributions of an element $e$ as a surrogate weight. In other words, if $f$ is the objective function, then the adaptations implicitly define the weight of $e$ to be $w(e)\triangleq f(A\cup\{e\})-f(A)$, where AA is the current solution. [^2] In some cases, AA is in fact some appropriate subset of the current solution, but this distinction is unimportant for the sake of this high-level discussion. However, applying this idea to the algorithm of Singer and Thiery [38] is problematic.
A significant bottleneck in the analysis of Singer and Thiery [38] occurs when there is a need to charge elements of the optimal solution to elements of the output solution that have nearly equal weight, but nevertheless belong to different weight classes. In this case, Singer and Thiery’s approximation guarantee degenerates to $k$, matching the guarantee of the greedy algorithm. This bottleneck affects only optimal elements whose weight lies at the end of the range of possible weights for their weight class. Thus, to overcome it Singer and Thiery apply a random shift to all weight classes, which allows them to bound the probability that any given optimal element has this undesirable property. In our setting, the surrogate “weight” of an element depends on the current solution AA being constructed by the algorithm, and thus is not independent of the random shift chosen; which means that it is no longer possible to bound the probability that an optimal element has a weight lying close to the end of the range of its weight class.
Our solution for the above hurdle is to introduce an additional auxiliary weight $u(e)$ for every element $e$ in the optimal solution $O$. The weights $u(e)$ depend only on $O$, rather than on the solution set AA, and thus are independent of the random bits of the algorithm. However, as these weights depend on the unknown set $O$, they are unavailable to the algorithm, which must continue to use the weights $w(e)$. This limitation of the algorithm is not important when $w(e)$ and $u(e)$ roughly agree, and thus, the output of the algorithm is good (in expectation) in this case. In order to obtain a result that holds in general, we further show that whenever the two weights disagree significantly, this discrepancy can be used to obtain an alternative bound on the value of the output of the algorithm.
Along the way, we also slightly modify the distribution of the random shift used by the algorithm to partition the elements into weight classes, which leads to some improvement also for linear objective functions compared to the approximation ratio of roughly $0.722k+O(1)$ obtained by [38]. More formally, we prove the following theorem. {restatable}theoremthmMonotoneResult For every $\varepsilon>0$, there exists an algorithm that runs in $\operatorname{Poly}(|E|,\varepsilon^{-1})$ time and guarantees an approximation ratio of $\frac{2k\ln 2}{1+\ln 2}+O(\sqrt{k})\leq 0.819k+O(\sqrt{k})$ for maximizing a non-negative monotone submodular function $f\colon 2^{E}\to{{\mathbb{R}}_{\geq 0}}$ under a matroid $k$-parity constraint. Moreover, if $f$ is linear, the approximation ratio of the algorithm improves to $k\ln 2+1-\ln 2+\varepsilon\leq 0.694k+0.307+\varepsilon$.
We note that the algorithm of [38] considers exchanges of size proportional to $k$ in its local search, which leads to a time complexity that is polynomial only when $k$ is constant. In contrast, the algorithm of Theorem 1.1 considers only exchanges of constant size, and therefore, runs in polynomial time regardless of the value of $k$.
We also consider the maximization of a non-monotone submodular function subject to a matroid $k$-parity constraint. Algorithms for maximizing monotone submodular functions usually still produce for a non-monotone submodular function $f$ a guarantee of the form $f(A)\geq\alpha\cdot f(A\cup O)$, where AA is the output set of the algorithm, $O$ is the optimal solution and $\alpha$ is positive number. Gupta, Roth, Schoenebeck and Talwar [18] described a general framework for translating such a guarantee into a real approximation ratio, and their framework was improved by subsequent works [33, 13]. Unfortunately, because of the use of our novel auxiliary weights, when our algorithm from Theorem 1.1 is analyzed in the context of a non-monotone submodular function $f$, it yields a guarantee of the more involved form $f(A)\geq\alpha\cdot f(A\cup O)-\beta\cdot f(O)$ for some $\beta<\alpha$. Nevertheless, we are able to adapt the version of the framework used by [13] to transform guarantees of our more involved form into standard approximation ratios, which yields the following theorem. {restatable}theoremthmNonMonotoneResult There exists an algorithm that runs in $\operatorname{Poly}(|E|)$ time and guarantees an approximation ratio of $\frac{2k\ln 2}{1+\ln 2}+O(k^{2/3})\leq 0.819k+O(k^{2/3})$ for maximizing a non-negative (not necessarily monotone) submodular function $f\colon 2^{E}\to{{\mathbb{R}}_{\geq 0}}$ under a matroid $k$-parity constraint.
1.2 Additional Related Work
When $k=2$, one can exactly optimize a linear function over a matroid $k$-intersection constraint via the classical algorithms of Edmonds [12] or Cunningham [10]. Similarly, a 22-set packing constraint is simply a standard matching constraint, and one can exactly optimize linear functions over it in polynomial time. Lawler [26] introduced the family of matroid parity constraints as a common generalization of both these families of constraints. Here, we are given a set $E$ of disjoint pairs of elements from $V$ and a matroid $\mathcal{M}$} on $V$, and the goal is to find a collection of pairs from $E$ whose union is independent for $\mathcal{M}$}. In contrast to the constraints they generalize, it is impossible to maximize even a cardinality objective under a matroid parity constraint in polynomial time in the setting in which the matroid is given as an independence oracle [31, 23]. Maximizing a cardinality objective under a matroid parity constraint is also known to be NP-hard (see, e.g., [37, §43.9]) and Lee, Sviridenko and Vondrák [29] showed that it admits a PTAS.
When $k\geq 3$, optimizing over both matroid $k$-intersection and $k$-set packing constraints is NP-hard as both these constraint families generalize $k$-dimensional matching constraints, and optimizing over the last family of constraints is NP-hard even for cardinality objectives [24]. All these types of constraints are generalized by the family of matroid $k$-parity constraints, and this family also generalizes the above-mentioned family of matroid parity constraints in the sense that matroid parity constraints are equivalent to matroid 22-parity constraints. The formal definition of matroid $k$-parity constraints can be found in Section 2, and we refer the reader to [29] for a more detailed discussion of the relationship between these and other families of constraints.
Finally, we note that improved results are known in several special cases. For matroid 22-intersection constraints, Huang and Sellier [20] give a $\frac{3}{2}$-approximation in the special case that the objective is a vertex cover function. For matroid $k$-parity constraints, there are several results for the special case that the matroid is linear and a representation is given. In this setting, when $k=2$ an exact polynomial time algorithm for cardinality objectives was given by Lovász [30], and Iwata and Kobayashi [21] later gave an exact polynomial time algorithm for maximizing linear objectives. For larger $k$ values, in the same setting, exact FPT algorithms are known for both cardinality and linear objectives, with exponential dependence on $k$ and the rank of the matroid or the number of blocks in the solution [2, 32]. In contrast, in this paper, we give approximation algorithms that require no assumptions on the structure of the underlying matroid and have a running time that is independent of $k$.
Paper Structure.
Section 2 describes the definitions and notation that we use, as well as some known results. Our main algorithm is then presented in Section 3, and Section 4 shows that this algorithm has the approximation guarantees stated by Theorem 1.1 for linear and monotone submodular functions. Finally, Section 5 explains how to use the main algorithm to prove also our result for non-monotone submodular functions (Theorem 1.1).
2 Preliminaries
In this section, we present preliminaries consisting of definitions, notations and known results used in this paper. The preliminaries are loosely grouped based on whether they are more related to the objective function or to the constraint of the problem we consider.
2.1 Constraint Preliminaries
Matroids.
A combinatorial constraint is given by a pair $(E,{\mathcal{I}})$, where $E$ is a ground set of elements and ${\mathcal{I}}\subseteq 2^{E}$ is the collection of all subsets of $E$ that are feasible according to the constraint. The constraint ${\mathcal{M}}=(E,{\mathcal{I}})$ is a matroid if $\mathcal{I}$} is non-empty and also obeys the following two properties.
  • Down-closedness: for every two sets $S\subseteq T\subseteq E$, if $T\in{\mathcal{I}}$ then also $S\in{\mathcal{I}}$.
  • Augmentation: for every two sets $S,T\in{\mathcal{I}}$, if $|S|<|T|$, then there exists an element $e\in T\setminus S$ such that $S\cup\{e\}\in{\mathcal{I}}$.
Like in the last definition, we often need in this paper to add or remove a single element from a set. For this purpose, given a set $S$ and an element $e$, it is convenient to use $S+e$ and SeS-e as shorthands denoting $S\cup\{e\}$ and $S\setminus\{e\}$, respectively. We also need to introduce some concepts from matroid theory. It is customary to refer to the sets in the collection $\mathcal{I}$} as the independent sets of $\mathcal{M}$}. Notice that the augmentation property of matroids implies that all the inclusion-wise maximal independent sets of $\mathcal{M}$} have the same size. This size is known as the rank of $\mathcal{M}$}, and the (inclusion-wise maximal) independent sets of this size are the bases of $\mathcal{M}$}. It is also useful to define for every set $S\subseteq E$ a value $\operatorname{rank}_{\mathcal{M}}(S)$ that is equal to the largest size of an independent subset $S^{\prime}$ of $S$. Notice that $\operatorname{rank}_{\mathcal{M}}(E)$ is equal to the rank of the matroid $\mathcal{M}$}.
Given a matroid ${\mathcal{M}}=(E,{\mathcal{I}})$, there are a few standard methods to produce a new matroid from it. In the following, we survey three of these methods.
  • Given a set $S\subseteq E$, the restriction of $\mathcal{M}$} to $S$ is a matroid over the ground set $S$ that is denoted by ${\mathcal{M}}|_{S}$. The independent sets of ${\mathcal{M}}|_{S}$ are all the independent sets $\mathcal{M}$} that are subsets of this ground set. More formally, ${\mathcal{M}}|_{S}=(S,2^{S}\cap{\mathcal{I}})$.
  • Consider a set $S\subseteq E$. Contracting $S$ in the matroid $\mathcal{M}$} produces a matroid over the ground set $E\setminus S$ that is denoted by ${\mathcal{M}}/S$. Let $B$ be a maximal independent subset of $S$. Then, a set $T\subseteq E\setminus S$ is independent in ${\mathcal{M}}/S$ if and only if $T\cup B\in{\mathcal{I}}$. [^3] There might be multiple maximal independent subsets of $S$, and therefore, multiple options for choosing $B$. It is known that ${\mathcal{M}}/S$ is the same matroid for every such choice of $B$.
  • Given an integer $0\leq r\leq\operatorname{rank}_{\mathcal{M}}(E)$, the truncation of $\mathcal{M}$} to the rank $r$ is a matroid over the same ground set denoted by $\operatorname{trunc}({\mathcal{M}},r)$. A set is independent in $\operatorname{trunc}({\mathcal{M}},r)$ if it is independent in $\mathcal{M}$} and its size is at most $r$. Notice that this implies, in particular, that the rank of $\operatorname{trunc}({\mathcal{M}},r)$ is $r$.
We end the introduction of matroids by presenting the following known result about them, which is known as the Greene–Magnanti Theorem [17].
Lemma 2.1.
Let $S$ and $T$ be bases of a matroid $\mathcal{M}$}. Then, for every partition $S_{1},S_{2},\dotsc,S_{k}$ of $S$, there exists a partition $T_{1},T_{2},\dotsc,T_{k}$ of $T$ such that $(S\setminus S_{i})\cup T_{i}$ is a base of $\mathcal{M}$} for every $i\in[k]$.
Matroid $k$-Parity.
A matroid $k$-parity constraint is given by a matroid ${\mathcal{M}}=(V,{\mathcal{I}}_{{\mathcal{M}}})$ together with a partition $E$ of the elements of $V$ into disjoint sets of size at most $k$. Borrowing terminology from hypergraphs, we refer to the elements of $E$ as edges and the elements of $V$ as vertices. The ground set of the matroid $k$-parity constraint is the set $E$ of edges and a subset $S\subseteq E$ is feasible according to the constraint if the union of the edges in $S$ is independent in $\mathcal{M}$}. More formally, if we denote by $v(e)$ the set of up to $k$ vertices of $V$ appearing in $e$, then the matroid $k$-parity constraint is the combinatorial constraint $(E,{\mathcal{I}})$, where
${\mathcal{I}}\triangleq\bigg\{S\subseteq E~\bigg|~\bigcup_{e\in S}v(e)\in{\mathcal{I}}_{\mathcal{M}}\bigg\}\kern 5.0pt.$
It is often useful to extend the function $v$ to sets of edges in the natural way, namely, $v(S)\triangleq\cup_{e\in S}\;v(e)$ for every set $S\subseteq E$. Note that every matroid $k$-parity constraint is down-closed; i.e., if $A\in{\mathcal{I}}$ then $B\in{\mathcal{I}}$ for all $B\subseteq A$. This follows immediately from the fact that the matroid $\mathcal{M}$} is down-closed by definition.
In Section 1, it was mentioned that matroid $k$-parity constraints generalize the intersection of $k$ matroid constraints. Let us now briefly explain why that is the case. Consider an arbitrary combinatorial constraint $(X,{\mathcal{I}})$ that can be obtained as the intersection of $k$ matroids, i.e., there exist $k$ matroids $(X,{\mathcal{I}}_{1}),(X,{\mathcal{I}}_{2}),\dotsc,\allowbreak(X,{\mathcal{I}}_{k})$ such that ${\mathcal{I}}=\cap_{i=1}^{k}{\mathcal{I}}_{k}$. We can represent this as a matroid $k$-parity constraint as follows. Define an edge for each element $x\in X$, and let the $k$ vertices of this edge be denoted by $(x,1),(x,2),\dotsc,(x,k)$. Intuitively, the element $(x,i)$ represents the element $x$ in the matroid $(X,{\mathcal{I}}_{i})$. Based on this intuition, we construct a matroid $\mathcal{M}$} over the resulting set $V$ of vertices by defining that a subset $S\subseteq V$ is independent in $\mathcal{M}$} if
$\{e\in E\mid(e,i)\in S\}\in{\mathcal{I}}_{i}\quad\forall\;i\in[k]\kern 5.0pt.$
One can verify that $\mathcal{M}$} is a matroid, and so $(E,{\mathcal{I}})$ is a matroid $k$-parity constraint given by $E$ and $\mathcal{M}$} that is equivalent to the original matroid intersection constraint.
We use in this paper the following exchange property of matroid $k$-parity constraints. This property was implicitly proved by [38], but we provide a different, and arguably simpler, proof of it.
Lemma $2.$
Consider a matroid $k$-parity constraint $(E,{\mathcal{I}})$, and let $A,B\in{\mathcal{I}}$ be two sets of edges that are feasible for this constraint. Then, there exists a collection of sets $\{N_{b}\subseteq A\mid b\in B\}$ such that
  1. 1.
    For every element $b\in B\cap A$, $N_{b}=\{b\}$; and for every element $b\in B\setminus A$, $N_{b}\subseteq A\setminus B$.
  2. 2.
    $A\cup\{b\in B:N_{b}=\emptyset\}\in{\mathcal{I}}$.
  3. 3.
    For every $a\in A$, $(A-a)\cup\{b\in B:N_{b}=\{a\}\}\in{\mathcal{I}}$.
  4. 4.
    For every $a\in A$, there are at most $k$ distinct $b\in B$ such that $a\in N_{b}$.
Proof.
Let ${\mathcal{M}}=(V,{\mathcal{I}}_{\mathcal{M}})$ and $E\subseteq 2^{V}$ be the matroid and partition defining the given matroid $k$-parity constraint. Suppose first that $A\cap B\neq\emptyset$. Then, we can define a new matroid $k$-parity constraint based on the matroid ${\mathcal{M}}/v(A\cap B)$ and the set of edges $E\setminus A\cap B$. Clearly, $A\setminus B$ and $B\setminus A$ are disjoint and feasible for this new matroid $k$-parity constraint. By applying the present lemma to these disjoint sets and this new constraint, we can obtain a collection of sets $\{N_{b}\subseteq A\setminus B\mid b\in B\setminus A\}$. Furthermore, adding to this collection the set $N_{b}=\{b\}$ for every $b\in A\cap B$ yields an extended collection of sets obeying all the properties required by the present lemma for the original sets AA and $B$ and the original matroid $k$-parity constraint. Thus, to prove the present lemma for general sets AA and $B$, it suffices to prove it for disjoint sets AA and $B$, which is our goal in the rest of this proof.
If $|v(A)|\leq|v(B)|$, then by repeated applications of the augmentation property of matroids, there is some set $C\subseteq v(B)\setminus v(A)$ with |C|=|v(B)||v(A)||C|=|v(B)|-|v(A)| such that $v(A)\cup C\in{\mathcal{I}}_{\mathcal{M}}$. Similarly, if $|v(A)|\geq|v(B)|$, then there is some set $C\subseteq v(A)\setminus v(B)$ with |C|=|v(A)||v(B)||C|=|v(A)|-|v(B)| such that $v(B)\cup C\in{\mathcal{I}}_{\mathcal{M}}$. Thus, in either case, there is a set $C$ such that $|v(A)\setminus C|=|v(B)\setminus C|$, $v(A)\cup C\in{\mathcal{I}}_{\mathcal{M}}$ and $v(B)\cup C\in{\mathcal{I}}_{\mathcal{M}}$.
Define now the matroid $\overline{{\mathcal{M}}}\triangleq\operatorname{trunc}(({\mathcal{M}}|_{v(A\cup B)})/C,|v(A)\setminus C|)$, and let $\overline{{\mathcal{I}}}\subseteq 2^{V\setminus C}$ be the collection of independent sets of $\overline{{\mathcal{M}}}$. For each element $e\in A\cup B$, let $\overline{v}(e)\triangleq v(e)\setminus C$, and similarly let $\overline{v}(S)\triangleq v(S)\setminus C=\bigcup_{e\in S}\overline{v}(e)$ for any $S\subseteq A\cup B$. Then, $\overline{v}(A)$ and $\overline{v}(B)$ are both bases of $\overline{{\mathcal{M}}}$—because they are independent in $\overline{{\mathcal{M}}}$ and their common size matches the rank of the truncation operation in the definition of $\overline{{\mathcal{M}}}$. The sets $\{\overline{v}(a)\mid a\in A\}$ form a partition of the base $\overline{v}(A)$, and thus, by Lemma 2.1, there exists a partition $\{\pi_{a}\subseteq\overline{v}(B)\mid a\in A\}$ of the base $\overline{v}(B)$ such that $|\pi_{a}|=|\overline{v}(a)|$ and $(\overline{v}(A)\setminus\overline{v}(a))\cup\pi_{a}\in\overline{{\mathcal{I}}}$ for every $a\in A$. This allows us to define $N_{b}\triangleq\{a\in A\mid\pi_{a}\cap\overline{v}(b)\neq\emptyset\}$ for each $b\in B$. In the following, we prove that this collection of sets obeys all the claims stated by the lemma.
Claim 1 follows immediately from our assumption that AA and $B$ are disjoint. Let us prove next claim 2. Define $S=\{b\in B\mid N_{b}=\emptyset\}$. For every $b\in S$, it holds by definition that $\pi_{a}\cap\overline{v}(b)=\emptyset$ for every $a\in A$. Since the sets $\pi_{a}$ are a partition of $\overline{v}(B)$ and $\overline{v}(b)\subseteq\overline{v}(B)$, we must have $\overline{v}(b)=\emptyset$ for every such $b\in S$. Thus,
$v(A\cup S)\subseteq\overline{v}(A\cup S)\cup C=\overline{v}(A)\cup C=v(A)\cup C\in{\mathcal{I}}\kern 5.0pt.$
For the proof of claim 3, fix $a\in A$ and let $S=\{b\in B\mid N_{b}=\{a\}\}$. For all $b\in S$, $\overline{v}(b)\subseteq\pi_{a}$ because otherwise $\overline{v}(b)$ would have intersected additional sets in the partition of $\overline{v}(B)$ beside $\pi_{a}$, which would have resulted in additional elements of AA appearing in $N_{b}$. Thus,
$\overline{v}((A-a)\cup S)=(\overline{v}(A)\setminus\overline{v}(a))\cup\overline{v}(S)\subseteq(\overline{v}(A)\setminus\overline{v}(a))\cup\pi_{a}\in\overline{{\mathcal{I}}}\kern 5.0pt,$
and therefore, $v((A-a)\cup S)\subseteq\overline{v}((A-a)\cup S)\cup C\in{\mathcal{I}}$.
Finally, for the proof of claim 4, fix $a\in A$ and recall that $v(b)$ and $v(b^{\prime})$ are disjoint for every two distinct elements $b,b^{\prime}\in B$. Therefore, $\pi_{a}\cap\overline{v}(b)\subseteq v(b)$ and $\pi_{a}\cap\overline{v}(b^{\prime})\subseteq v(b^{\prime})$ must also be disjoint for every two such elements. Thus,
$|\{b\in B\mid a\in N_{b}\}|=|\{b\in B\mid\pi_{a}\cap\overline{v}(b)\neq\emptyset\}|\leq\sum_{b\in B}|\pi_{a}\cap\overline{v}(b)|=|\pi_{a}|=|\overline{v}(a)|\leq|v(a)|\leq k\kern 5.0pt.\qed$
Independence/Feasibility Oracle.
We conclude the the discussion of combinatorial constraints by noticing that in most of their applications it is useful to have algorithms whose time complexities are polynomial in the size of the ground set $E$ rather than the size of the collection $\mathcal{I}$} of feasible sets, which might be exponential in $|E|$. Therefore, algorithms that operate on such constraints usually assume that $\mathcal{I}$} is not part of the input, but can be accessed through an oracle, known as independence or feasibility oracle, that takes as input a set $S\subseteq E$ and answers whether $S\in{\mathcal{I}}$. The algorithms in this paper implicitly follow this standard assumption.
2.2 Objective Function Preliminaries
A set function $f\colon 2^{E}\to{\mathbb{R}}$ is a function that assigns a value to every subset of a given ground set $E$. Given a set $S$ and an element $e$, we denote by $f(e\mid S)\triangleq f(S+e)-f(S)$ the marginal contribution of $e$ to $S$. Similarly, given an additional set $T$, we denote by $f(T\mid S)\triangleq f(S\cup T)-f(S)$ the marginal contribution of $T$ to $S$. The set function $f$ is called monotone if $f(e\mid S)\geq 0$ for every set $S\subseteq E$ and element $e\in E$. The set function $f$ is called submodular if $f(e\mid S)\geq f(e\mid T)$ for every two sets $S\subseteq T$ and element $e\in E\setminus T$. In other words, $f$ is submodular if the marginal contribution of an element $e$ to a set $S$ can only diminish when additional elements are added to the set $S$. An alternative, equivalent, characterization of submodularity states that $f$ is submodular if and only if $f(S)+f(T)\geq f(S\cup T)+f(S\cap T)$ for every two sets $S,T\subseteq E$.
In the case that $f(S)+f(T)=f(S\cup T)+f(S\cap T)$ for all $S,T\subseteq E$, we say that $f$ is modular. Equivalently, a modular set function can be defined as a function of the form $f(S)=w_{0}+\sum_{e\in S}w_{e}$, where $w_{0}$ and $\{w_{e}\mid e\in E\}$ are arbitrary real values. In the special case that $w_{0}=0$ (and thus, $f(\emptyset)=0$), we say that $f$ is linear.
Our algorithms optimize a submodular set function $f$. However, as in the case of combinatorial constraints, it often does not make sense to assume that the algorithm has an explicit description of $f$ because this description can be exponential in the size of the ground set $E$. Instead, our algorithms implicitly make the standard assumption that $f$ is accessed through a value oracle that given a set $S\subseteq E$ returns $f(S)$.
3 Main Algorithm
In this section, we present our main algorithm, given below as Algorithm 1. In Section 4, we prove that Algorithm 1 provides the approximation guarantees for linear and monotone submodular functions given in Theorem $1.$ Algorithm 1 is also a central component in the proof of Theorem 1.1 for non-monotone submodular objectives. As stated, Algorithm 1 is not necessarily a polynomial time algorithm. However, in Appendix A we describe a polynomial time implementation of it.
Algorithm 1 gets as input a ground set $E$, a submodular function $f\colon 2^{E}\to{\mathbb{R}}$ obeying $f(\emptyset)\geq 0$, the collection $\mathcal{I}$} of the feasible sets of some matroid $k$-parity constraint over the ground set $E$ and an error parameter $\varepsilon\in(0,1)$. It then iteratively constructs sets $A_{1},A_{2},\dotsc$ whose union becomes the algorithm’s output. For ease of notation, Algorithm 1 uses $A_{\leq i}$ to denote the union $A_{1}\cup A_{2}\cup\dotso\cup A_{i}$. When thinking in terms of the intuition given in Section 1, $A_{1}$ is the part of the solution of the algorithm chosen from the highest “weight class”, $A_{2}$ is the part of the solution of the algorithm chosen from the second highest “weight class”, and so on. Accordingly, Algorithm 1 adds to the set $A_{i}$ only elements with large enough marginal values. More specifically, Algorithm 1 defines a series of exponentially decreasing thresholds $m_{i}$, and adds to $A_{i}$ only elements whose marginal contribution is at least $m_{i}$.
To understand the details of the construction of the sets $A_{i}$ by Algorithm 1, we need the following definition.
Definition 3.1 ($(\theta,\varepsilon)$-improvement).
Let $\theta>0$ and $\varepsilon\in(0,1)$ be two given values, and $A\in{\mathcal{I}}$ be a feasible solution. We say that a pair of sets $S\subseteq E\setminus A$ and $N\subseteq A$ is a $(\theta,\varepsilon)$-improvement for AA if $(A\cup S)\setminus N\in{\mathcal{I}}$ and one of the following holds:
  1. 1.
    $S=\{x\},N=\emptyset$, and $f(x\mid A)\geq\theta$.
  2. 2.
    $S=\{x\},N=\{y\}$, $f(x\mid A)\geq\theta$, and $f(A+x-y)\geq f(A)+\varepsilon\theta$.
  3. 3.
    $|S|=2,N=\{y\}$ and $f(x_{1}\mid A)\geq\theta$, $f(x_{2}\mid A+x_{1})\geq\theta$ for at least one of the two possible ways to assign the labels $x_{1},x_{2}$ to the elements of $S$.
At the beginning of iteration $i$ of Algorithm 1, the sets $A_{1},A_{2},\dotsc,A_{i-1}$ are already finalized. Algorithm 1 then tries to construct a set $A_{i}$ that is a good addition to these sets. It does so by initializing $A_{i}$ to be the empty set, and then improving $A_{\leq i-1}\cup A_{i}$ via a series of $(m_{i},\varepsilon)$-improvements. Notice that, by definition, the elements added by these improvements to $A_{i}$ have marginals of at least $m_{i}$, as promised above. Since the sets $A_{1},A_{2},\dotsc,A_{i-1}$ are finalized before the $i$-th iteration of Algorithm 1, and must not be modified at this iteration, the algorithm considers only $m_{i}$-improvements whose set NN of removed elements is a subset of the current set $A_{i}$. Whenever an improvement of this kind is found, it is applied by setting $A_{i}\leftarrow(A_{i}\cup S)\setminus N$, and once no more such $m_{i}$-improvements can be found, the $i$-th iteration of Algorithm 1 terminates.
1Let $W\leftarrow\max_{e\in E}f(e\mid\emptyset)$.
2 Let $\alpha$ be a uniformly at random value from $(0,1]$, and let $\tau\leftarrow 2^{\alpha}$.
3 Define $m_{i}\triangleq W\tau 2^{-i}$ for every integer $i\geq 0$.
4
51ex
6Let $i\leftarrow 0$ and $A_{\leq 0}\leftarrow\emptyset$.
7 while there exists $e\in E\setminus A_{\leq i}$ such that $f(e\mid A_{\leq i})>0$ and $A_{\leq i}+e\in{\mathcal{I}}$ do
8   Update $i\leftarrow i+1$.
9   Let $A_{i}\leftarrow\emptyset$.
10 while there is some $(m_{i},\varepsilon)$-improvement $(S,N)$ for $A_{\leq i-1}\cup A_{i}$ with $N\subseteq A_{i}$ do
11    $A_{i}\leftarrow(A_{i}\setminus N)\cup S$.
12
13  Let $A_{\leq i}\leftarrow A_{\leq i-1}\cup A_{i}$.
14
return $A_{\leq i}$.
Algorithm 1 Greedy/Local-Search Hybrid Algorithm $(E,f,{\mathcal{I}},\varepsilon)$
In Appendix A, we show that Algorithm 1 can in fact be simulated in polynomial time. However, in order to derive the approximation guarantees of Theorem 1.1, it is sufficient to prove here the weaker claim that Algorithm 1 terminates. In each iteration $i$, Algorithm 1 repeatedly searches for and applies $(m_{i},\varepsilon)$-improvements for some current solution AA. Each such improvement either increases $|A|$, or leaves $|A|$ unchanged and increases $f(A)$. Since there are only finitely many possible distinct values for $|A|$ and $f(A)$, it follows that each iteration of Algorithm 1 must terminate. The following observation implies that Algorithm 1 can perform at most finitely many iterations before terminating.
Observation 3.1.
For any input to Algorithm 1, there is an integer $L\geq 0$ such that $f(e\mid A_{\leq L})\leq 0$ for every $e\in E\setminus A_{\leq L}$ with $A_{\leq L}+e\in{\mathcal{I}}$, and furthermore, if $L\neq 0$, then $A_{L}\neq\emptyset$.
Proof.
First, we note that any iteration $i$ in which at least one $(m_{i},\varepsilon)$-improvement is found, must end with $|A_{i}|>0$. This follows directly from the structure of $(\theta,\varepsilon)$-improvements, which never decrease the size of the set AA that is being improved. Since the sets $A_{i}$ produced across different iterations are disjoint by construction, there can be at most $|E|$ iterations in which the algorithm finds at least one $(\theta,\varepsilon)$-improvement. Let $L$ be the index of the last such iteration, or 0 if there are no such iterations. Suppose, for the sake of contradiction, that there is some $e\in E\setminus A_{\leq L}$ with $A_{\leq L}+e\in{\mathcal{I}}$ and $f(e\mid A_{\leq L})>0$. Since no $(\theta,\varepsilon)$-improvements are found after iteration $L$, we must have $A_{\leq i}=A_{\leq L}$ for all $i>L$. But, for sufficiently large $i\geq 1+\log_{2}W-\log_{2}f(e\mid A_{\leq L})$, we then have $e\in E\setminus A_{\leq i}$ with $f(e\mid A_{\leq i})=f(e\mid A_{\leq L})\geq m_{i}$ and $A_{\leq i}+e=A_{\leq L}+e\in{\mathcal{I}}$. Thus, $(\{e\},\emptyset)$ would be a valid $(m_{i},\varepsilon)$-improvement in iteration $i>L$, contradicting the fact that no such improvements are found by Algorithm 1. ∎
Observation 3.1 implies that Algorithm 1 terminates after $L$ iterations. Therefore, in the rest of our analysis we use $A_{\leq L}$ to denote the output set of Algorithm 1. Note that the condition on Line 1 of the algorithm implies that the set $A_{\leq i}=A_{\leq i-1}\cup A_{i}$ finalized at the end of each iteration $i$ is locally optimal with respect to $(m_{i},\varepsilon)$-improvements. That is, for every $i\in[L]$, there is no $(m_{i},\varepsilon)$ improvement $(S,N)$ for $A_{\leq i}$ with $N\subseteq A_{i}$. In particular, considering $(m_{i},\varepsilon)$-improvements of the first kind gives the following observation.
Observation 3.2.
For every $i\in[L]$, the inequality $f(e\mid A_{\leq i})<m_{i}$ holds for every $e\in E\setminus A_{\leq i}$ that obeys $A_{\leq i}+e\in{\mathcal{I}}$ at the end of iteration ii.
Proof.
Suppose $A_{\leq i}+e\in{\mathcal{I}}$ for some $e\in E\setminus A_{\leq i}$ at the end of iteration ii. If $f(e\mid A_{\leq i})\geq m_{i}$, then $(\{e\},\emptyset)$ would be a valid $(m_{i},\varepsilon)$-improvement, contradicting the fact that the loop on Line 1 of Algorithm 1 terminated. ∎
4 Analysis of the Main Algorithm
In this section, we bound the approximation guarantee of Algorithm 1 when $f\colon 2^{E}\to{\mathbb{R}}$ is a general submodular function with $f(\emptyset)\geq 0$. We then use this bound to prove Theorem $1.$ Fix a matroid $k$-parity instance $(E,{\mathcal{I}})$. In the following, we consider an execution of Algorithm 1 on this matroid $k$-parity and the above function $f$. Recall that $A_{\leq L}$ is the output of this algorithm, and thus, our objective is to lower bound $f(A_{\leq L})$ in terms of $f(O)$, where $O$ is an optimal solution for our given instance (i.e., a set in $\mathcal{I}$} maximizing $f$). Without loss of generality, we may restrict our analysis to optimal solutions that satisfy the following property.
Definition 4.1.
A feasible set $O$ is strictly down-monotone (with respect to $f$) if $f(x\mid O-x)>0$ for every $x\in O$.
To see that such a restriction is without loss of generality, note that as long as $f(x\mid O-x)\leq 0$ for some element $x\in O$, removing this element from $O$ gives a new feasible solution that is smaller, but has at least the same value as $O$, and thus, is also optimal. Hence, any minimal size optimal solution is strictly down-monotone.
Our analysis below does not use any property of $O$ beside its strict down-monotonicity. In particular, it does not use the optimality of $O$, and therefore, applies also when $O$ is a strictly down-monotone set that is not necessarily optimal. This property of the analysis is used in Section 5. The analysis consists of a few steps. First, in Section 4.1, we partition the elements of $O$ into sets $\{O_{i}\mid i\in[L]\}$. Intuitively, each $O_{i}$ is a subset of $O$ whose contribution is bounded in terms of the contribution of the elements of $A_{i}$, i.e., the elements selected in iteration $i$ of Algorithm 1. We quantify these contributions in Section 4.2 by defining weights $w$ and $\bar{w}$} for the elements of $A_{\leq L}$ and $O$, respectively. Section 4.3 then classifies the elements of $O$ into three different types of elements, and Section 4.5 separately bounds the total weight of each type in terms of the weight of $A_{\leq L}$ via appropriate charging arguments.
In the linear case, the weights $w$ and $\bar{w}$} of elements are identical and independent of the sets Algorithm 1 constructs. This makes our charging scheme essentially the same as the one used in [38]. However, in the submodular case, the weights depend on the final set $A_{\leq L}$ returned by the algorithm. In order to analyze the effect of the random shift $\tau$ selected by Algorithm 1, we must thus introduce, in Section 4.4, a set of auxiliary weights $u$ quantifying the value of $O$ in a fashion that is independent of $A_{\leq L}$. Our analysis then explicitly considers how these weights $u$ differ from $\bar{w}$}. Specifically, we show that if there is a large discrepancy between $u$ and $\bar{w}$}, then we can handle the resulting loss in the analysis by balancing with an alternative lower bound on $f(A_{\leq L})$ that gains from such a discrepancy.
4.1 Partitioning the Elements of $O$
This section describes a partition of $O$ into disjoint sets $O_{1},O_{2},\dotsc,O_{L}$. Analogously to the sets $A_{\leq i}$, for every integer $0\leq i\leq L$, we use $O_{\leq i}$ to denote the union $\cup_{j=1}^{i}O_{j}$. Intuitively, each set $O_{i}$ represents elements from $O$ whose value can be bounded in terms of the elements selected as $A_{i}$ in iteration $i$ of Algorithm 1, and thus, need not be considered when later iterations of this algorithm are analyzed. In line with this intuition, we prove below that the elements of $O\setminus O_{\leq i}$, whose value needs to be bounded in terms of elements selected by later iterations, can all be feasibly added to the solution $A_{\leq i}$ that Algorithm 1 has after iteration ii.
We can now formally define the partition of $O$. Recall that $O_{\leq 0}=\emptyset$ by definition. For every $i\in[L]$, the set $O_{i}$ is given by the recursive formula
$O_{i}\triangleq\{o\in O\setminus O_{\leq i-1}\mid N^{(i)}_{o}\neq\emptyset\}\kern 5.0pt$,
where $\{N_{o}^{(i)}\subseteq A_{\leq i}\mid o\in A_{i-1}\cup(O\setminus O_{\leq{i-1}})\}$ is the collection of sets obtained by applying Lemma 2.2 with the sets $A_{\leq i}$ and $A_{\leq i-1}\cup(O\setminus O_{\leq i-1})$ as the sets AA and $B$ of the lemma, respectively. Recall that applying Lemma 2.2 with these sets requires them to be feasible. The feasibility of $A_{\leq i}$ is clear. The feasibility of $A_{\leq i-1}\cup(O\setminus O_{\leq i-1})$ follows from the next observation.
Observation 4.1.
For every integer $0\leq i\leq L$, $A_{\leq i}\cup(O\setminus O_{\leq i})$ is feasible and $A_{\leq i}\cap(O\setminus O_{\leq i})=\emptyset$.
Proof.
We prove the observation by induction on ii. First, note that the observation is trivial for $i=0$ because $A_{\leq 0}\cup(O\setminus O_{\leq 0})=O$ and $A_{\leq 0}=\emptyset$. Assume now that the observation holds for every integer $0\leq i^{\prime}<i$, and let us prove it for ii.
The induction hypothesis implies that $A_{\leq i^{\prime}}\cup(O\setminus O_{\leq i^{\prime}})$ is feasible for every $0\leq i^{\prime}<i$, and therefore, the consruction of $O_{i}$ via the above process is well-defined. The first part of the observation now follows from claim 2 of Lemma 2.2 because $O\setminus O_{\leq i}=\{o\in O\setminus O_{\leq i-1}\mid N^{(i)}_{o}$$\emptyset\}$, and the second part of the observation holds because claim 1 of Lemma 2.2 guarantees that every element $o$ that belongs both to $A_{\leq i}$ and $O\setminus O_{\leq i-1}$ has $N^{(i)}_{o}=\{o\}$, and therefore, belongs to $O_{i}$. ∎
4.2 Defining Weights for the Charging Argument
We now define the weights we use to quantify the values of elements in $O$ and $A_{\leq L}$. An element’s weight is based on the incremental value it adds to a given set, and in the linear case it corresponds exactly to the weight of this element in the definition the objective function $f$ (as long as this weight is non-negative). In order to define the incremental values, we order the elements of $A_{\leq L}$ as $a_{1},a_{2},\dotsc,a_{|A_{\leq L}|}$, in the order in which they were last added by Algorithm 1 to any of the sets $A_{1},A_{2},\dotsc,A_{L}$. Applying $(\theta,\varepsilon)$-improvements of the third kind involves adding 22 elements to $A_{i}$ at the same time, which creates an ambiguity. To resolve the ambiguity, recall that in such improvements, Definition 3.1 specifies that there must be a way to assign the two added elements the labels $x_{1}$ and $x_{2}$ so that $f(x_{1}\mid A)\geq\theta$ and $f(x_{2}\mid A+x_{1})\geq\theta$, where AA is the set improved by the $(\theta,\varepsilon)$-improvement. In our ordering of $A_{\leq L}$, we suppose that the element corresponding to $x_{1}$ is added to $A_{i}$ immediately before the element corresponding to $x_{2}$. [^4] If both ways to label the added elements obey the above inequalities, then we arbitrary select one of them. Given our ordering $a_{1},\dotsc,a_{|A_{\leq L}|}$, we define the weight of each element $a_{j}\in A_{\leq L}$ as
$w(a_{j})\triangleq f(a_{j}\mid\{a_{1},a_{2},\dotsc,a_{j-1}\})\kern 5.0pt.$
We proceed similarly for elements in $O$. Let us denote the elements of $O$ by $o_{1},o_{2}\dotsc,o_{|O|}$ in an arbitrary order. [^5] While the order $o_{1},o_{2}\dotsc,o_{|O|}$ is arbitrary, it is essentially that this order is not chosen in a way that depends on the random bits of Algorithm 1. Then, for any $o_{j}\in O$, we define
${\bar{w}}(o_{j})\triangleq\max\{0,f(o_{i}\mid(A_{\leq L}-o_{j})\cup\{o_{1},o_{2},\dotsc,o_{j-1}\})\}\kern 5.0pt.$
Note that we ensure that the $\bar{w}$} weights are always non-negative. Let us briefly provide some intuition for the second expression in the maximum. For any $o_{j}\in O\setminus A_{\leq L}$, this term is equal to $f(o_{j}\mid A_{\leq L}\cup\{o_{1},\dotsc,o_{j-1}\})$, and thus, the weights $\bar{w}$} capture the incremental values obtained when adding all such elements sequentially to the set $A_{\leq L}$. For any element $o_{j}\in O\cap A_{\leq L}$, we could define ${\bar{w}}(o_{j})=0$, since it provides no such incremental value. However, it is useful in the linear case to ensure that all the weights we are defining are equal to the natural notion of an element’s weight (i.e., $f(\{x\})$) whenever this weight is non-negative. The chosen expression ensures that this is the case even for $x\in O\cap A_{\leq L}$.
It is convenient to extend the above-defined weights to sets. We do this in the natural way, i.e., for any set $T\subseteq A_{\leq L}$, we define $w(T)\triangleq\sum_{e\in T}w(e)$, and for any set $T\subseteq O$, we define ${\bar{w}}(T)\triangleq\sum_{e\in T}{\bar{w}}(e)$. The following lemma proves some basic properties of the weights $w$ and $\bar{w}$}. The first three claims of the lemma show that an element’s weight must fall into a particular range defined by the values $m_{i}$, depending on which (if any) set $A_{i}$ or $O_{i}$ this element appears in. The final claim relates the two weights $w$ and $\bar{w}$} given to elements that appear in both $A_{\leq L}$ and $O$.
Lemma 4.2.
The weights $w$ and $\bar{w}$} satisfy the following.
  1. 1.
    For every $i\in[L]$ and element $a\in A_{i}$, $0<m_{i}\leq w(a)\leq m_{i-1}$.
  2. 2.
    For every $i\in[L]$ and element $o\in O_{i}$, ${\bar{w}}(o)\leq m_{i-1}$.
  3. 3.
    For all $o\in O\setminus O_{\leq L}$, ${\bar{w}}(o)=0$.
  4. 4.
    For every element $o\in O\cap A_{\leq L}$, ${\bar{w}}(o)\leq w(o)$.
Proof.
We prove each claim in turn.
Claim 1.
Consider the moment that $a\in A_{i}$ was last added to $A_{i}$ by Algorithm 1, and let $\bar{A}_{i}$ denote the value of this set just before the element $a$ was added by some $(m_{i},\varepsilon)$-improvement (if the $(m_{i},\varepsilon)$-improvement adds two elements, and $a$ is the one labeled $x_{2}$ in this improvement, then $\bar{A}_{i}$ is assumed to already contain the element labeled as $x_{1}$ by the improvement). Since $a$ is added by an $(m_{i},\varepsilon)$-improvement for the set $\bar{A}_{i}\cup A_{\leq i-1}$,
$m_{i}\leq f(a\mid\bar{A}_{i}\cup A_{\leq i-1})\leq w(a)\kern 5.0pt,$
where the second inequality follows from the submodularity of $f$ because all the elements of $A_{\leq L}$ that were last added to the sets $A_{1},A_{2},\dotsc,A_{L}$ before the last time in which $a$ was added to $A_{i}$ must also belong to $\bar{A}_{i}\cup A_{\leq i-1}$.
For the upper bound of the claim, first suppose that $a\in A_{1}$. Then,
$w(a)\leq f(a\mid\emptyset)\leq W\leq W\tau=m_{0}$
by submodularity and the definition of $W$. In the general case, suppose $a\in A_{i}$ for some $i>1$. Then, $a\in E\setminus A_{\leq i-1}$ with $\bar{A}_{i}\cup A_{\leq i-1}+a\in{\mathcal{I}}$. Since $\mathcal{I}$} is down-closed, $A_{\leq i-1}+a\in{\mathcal{I}}$. Observation 3.2 then implies that $m_{i-1}>f(a\mid A_{\leq i-1})\geq w(a)$, where the final inequality holds by submodularity since all the elements of $A_{\leq i-1}$ were added to the algorithm’s solution before $a$ was added to $A_{i}$.
Claim 2.
We have $o\in O_{i}\subseteq O\setminus O_{\leq i-1}$. By Observation 4.1, $(O\setminus O_{\leq i-1})\cap A_{\leq i-1}=\emptyset$ and $(O\setminus O_{\leq i-1})\cup A_{\leq i-1}\in{\mathcal{I}}$; and thus, $o\in E\setminus A_{\leq i-1}$ with $A_{\leq i-1}+o\in{\mathcal{I}}$. If $i>1$, then by Observation 3.2, we must then have $f(o\mid A_{\leq i-1})<m_{i-1}$. If $i=1$, we get the slightly weaker inequality $f(o\mid A_{\leq i-1})\leq f(o\mid\varnothing)\leq W\leq m_{0}=m_{i-1}$ by the submodularity of $f$. In either case, since $m_{i-1}$ is positive, $\max\{0,f(o\mid A_{\leq i-1})\}\leq m_{i-1}$. Suppose that $o$ is the $j$-th element $o_{j}$ in our ordering of $O$. Then, by submodularity,
${\bar{w}}(o)={\bar{w}}(o_{j})={}$ $\max\{0,f(o_{j}\mid(A_{\leq L}-o_{j})\cup\{o_{1},o_{2},\dotsc,o_{j-1}\})\}$
$\leq{}$ $\max\{0,f(o_{j}\mid A_{\leq{i-1}})\}\leq m_{i-1}\kern 5.0pt.$
Claim 3.
By Observation 4.1, $O\setminus O_{\leq L}$ and $A_{\leq L}$ are disjoint and $(O\setminus O_{\leq L})\cup A_{\leq L}\in{\mathcal{I}}$. Thus, for every $o_{j}\in O\setminus O_{\leq L}$ we have $o_{j}\in E\setminus A_{\leq L}$ and $A_{\leq L}+o_{j}\in{\mathcal{I}}$. Observation 3.1 then implies that $f(o_{j}\mid A_{\leq L})\leq 0$. Thus, by the submodularity of $f$,
${\bar{w}}(o_{j})=\max\{0,f(o_{j}\mid(A_{\leq L}-o_{j})\cup\{o_{1},o_{2},\dotsc,o_{j-1}\})\}\leq\max\{0,f(o_{j}\mid A_{\leq L})\}=0\kern 5.0pt.$
Claim 4
Suppose $o=o_{j}\in A_{\leq L}\cap O$. Then, we must have $o=a_{j^{\prime}}$ for some $a_{j^{\prime}}\in A_{\leq L}$. The first claim of the present lemma yields $w(a_{j^{\prime}})>0$ since $a_{j^{\prime}}$ must appear in some set $A_{i}$. Thus, by submodularity,
$w(o)=w(a_{j^{\prime}})=\max\{0,w(a_{j^{\prime}})\}$ $=\max\{0,f(a_{j^{\prime}}\mid\{a_{1},\dotsc,a_{j^{\prime}-1}\})\}$
$\geq\max\{0,f(a_{j^{\prime}}\mid(A_{\leq L}-a_{j^{\prime}})\cup\{o_{1},\dotsc,o_{j-1}\})\}$
$=\max\{0,f(o_{j}\mid(A_{\leq L}-o_{j})\cup\{o_{1},\dotsc,o_{j-1}\})\}={\bar{w}}(o)\kern 5.0pt.\qed$
4.3 Classifying the Elements of $O$ for Charging
In our charging scheme, we consider three different types of elements of $O$. The first is the elements that belong to some $O_{i}$, have weights approximately equal to the weights of the elements in $A_{i}$ (up to a factor of $m_{i-1}/m_{i}=2$) and can charge to only a single element of $A_{i}$. We denote below by $O^{(s)}$ the set of these elements, and they are intuitively treated using an argument borrowed from the unweighted $k$-matroid parity problem. The next type is the elements of $O\setminus O_{\leq L}$, which are “left over” at the end of Algorithm 1. The contributions of these elements are small, and can be essentially ignored. The final type is all remaining elements, i.e., $O_{\leq L}\setminus O^{(s)}$.
To formally define the set $O^{(s)}$ of the first type of elements of $O$, we define a set $N_{o}$ for every element $o\in O_{\leq L}$ as follows. For any $o\in O_{\leq L}$, let $i$ be the single integer in $[L]$ for which $o\in O_{i}$, and define $N_{o}\triangleq N^{(i)}_{o}$ (recall that $N^{(i)}_{o}$ is defined in Section 4.1). Intuitively, one can think of $N_{o}$ as the set of elements of $A_{\leq L}$ that are responsible for the inclusion of $o$ in $O_{i}$. We then define for every $i\in[L]$ a subset $O_{i}^{(s)}\subseteq O_{i}$ given by
$O_{i}^{(s)}$ $\triangleq\{o\in O_{i}\mid{\bar{w}}(o)>m_{i}\text{ and }|N_{o}|=1\}\kern 5.0pt.$
Then, we can define $O^{(s)}\triangleq\bigcup_{i=1}^{L}O_{i}^{(s)}$. Notice that this definition is consistent with the above intuition in the following senses. First, each element $o\in O^{(s)}$ has indeed only a single element in $N_{o}$ to “blame” for their inclusion in $O_{i}$, and thus, only a single element to charge to. Second, the weight of $o$ is comparable to the weight of this single element because $N_{o}\subseteq A_{i}$, which implies by claim 1 of Lemma 4.2 that the weight of every element in $N_{o}$ is in the range $m_{i},m_{i-1}]$.
The following lemma states some properties of the sets defined above. The first two claims of the lemma show that every element $o\in O^{(s)}$ “blames” a different element of $A_{\leq L}$ for its inclusion in $O_{\leq L}$, and furthermore, they show that this element is almost as valuable as $o$. Claim 3 of the lemma shows that every element of $A_{\leq L}$ is “blamed” by at most $k$ elements of $O_{\leq L}$.
Lemma 4.3.
The above defined sets obey the following.
  1. 1.
    For every $o\in O^{(s)}$, ${\bar{w}}(o)\leq(1+\varepsilon)\cdot w(N_{o})$.
  2. 2.
    For every two elements $o,o^{\prime}\in O^{(s)}$, $N_{o}\cap N_{o^{\prime}}=\emptyset$.
  3. 3.
    Every $a\in A_{\leq L}$ appears in $N_{o}$ for at most $k$ distinct $o\in O_{\leq L}$.
Proof.
First, we show that all $o\in O_{i}^{(s)}\setminus A_{\leq i}$ must have relatively large marginal contribution with respect to the set $A_{\leq i}$. By definition, any $o_{j}\in O_{i}^{(s)}$ has $0<m_{i}<{\bar{w}}(o_{j})=f(o_{j}\mid(A_{\leq L}-o_{j})\cup\{o_{1},\dotsc,o_{j-1}\})$. If $o_{j}\not\in A_{\leq i}$, then $(A_{\leq L}-o_{j})\cup\{o_{1},\dotsc,o_{j-1}\})\supseteq A_{\leq i}\cup O^{\prime}$ for any $O^{\prime}\subseteq\{o_{1},\dotsc,o_{j-1}\}$, and thus, by submodularity,
$m_{i}<{\bar{w}}(o_{j})\leq f(o_{j}\mid A_{\leq i})\,$, (1)
and also
$m_{i}<{\bar{w}}(o_{j})\leq f(o_{j}\mid A_{\leq i}+o_{j^{\prime}})\,$, (2)
for any $j^{\prime}<j$. Intuitively, these inequalities show that any such $o_{j}$ has marginal contribution large enough to be part of an $(m_{i},\varepsilon)$-exchange at the end of the iteration ii. Using these observations, we now prove each claim of the lemma in turn.
Claim 1.
Consider any $o\in O^{(s)}$ and let $i\in[L]$ be the (unique) value such that $o\in O^{(s)}_{i}$. Then, since $o\in O_{i}\subseteq O\setminus O_{\leq i-1}$, we must have $o\not\in A_{\leq i-1}$ by Observation 4.1. If $o\in A_{i}\cap O_{i}^{(s)}\subseteq A_{i}\cap O_{i}$, then by claim 1 of Lemma 2.2, we are guaranteed that $N_{o}=\{o\}$, which implies $w(o)=w(N_{0})$, and claim 4 of Lemma 4.2 further gives ${\bar{w}}(o)\leq w(o)$. Thus, ${\bar{w}}(o)\leq(1+\varepsilon)\cdot{\bar{w}}(o)\leq(1+\varepsilon)\cdot w(o)=(1+\varepsilon)\cdot w(N_{o})$, as required.
Now, suppose that $o\in O_{i}^{(s)}\setminus A_{i}$. Then, $o\not\in A_{\leq i}$, and $N_{o}=N_{o}^{(i)}=\{a\}$ for some element $a\in A_{i}$. By the properties of $N_{o}^{(i)}$ (specifically, Lemma 2.2, claim 3), we have $A_{\leq i}-a+o\in{\mathcal{I}}$. As shown above in (1), $f(o\mid A_{\leq i})>m_{i}$. Thus, $(\{o\},\{a\})$ satisfies all the properties required of an $(m_{i},\varepsilon)$ improvement of the second type, except for the inequality $f(A_{\leq i}-a+o)-f(A_{\leq i})\geq\varepsilon m_{i}$. Since Algorithm 1 terminated iteration $i$, and finalized the set $A_{\leq i}$, without applying this improvement, it must have been the case that this inequality was not satisfied, i.e.,
$\varepsilon m_{i}>{}$ $f(A_{\leq i}-a+o)-f(A_{\leq i})=f(o\mid A_{\leq i}-a)-f(a\mid A_{\leq i}-a)$
$\geq{}$ $f(o\mid A_{\leq i})-f(a\mid A_{\leq i}-a)\geq{\bar{w}}(o)-f(a\mid A_{\leq i}-a)\geq{\bar{w}}(o)-w(a)\kern 5.0pt,$
where the second inequality follows from submodularity, the third inequality holds by Inequality (1), and the last inequality follows from submodularity and the definition of $w$. Rearranging, we have
${\bar{w}}(o)\leq w(a)+\varepsilon m_{i}\leq(1+\varepsilon)\cdot w(a)=(1+\varepsilon)\cdot w(N_{o})\kern 5.0pt,$
where the second inequality follows from claim 1 of Lemma 4.2 since $a\in A_{i}$.
Claim 2.
Consider any distinct $o_{j},o_{j^{\prime}}\in O^{(s)}$, where $j^{\prime}<j$. We show below that if $N_{o_{j^{\prime}}}\cap N_{o_{j}}\neq\emptyset$, then an $(m_{i},\varepsilon)$-improvement of the third form exist for some $A_{\leq i}$, contradicting the fact that iteration $i$ of Algorithm 1 terminated and finalized the set $A_{\leq i}$ without applying this improvement.
To this end, suppose that $N_{o_{j^{\prime}}}\cap N_{o_{j}}\neq\emptyset$. Since both $o_{j^{\prime}}$ and $o_{j}$ are in $O^{(s)}$, we must have $N_{o_{j^{\prime}}}=N_{o_{j}}=\{a\}$ for some single element $a$, and this element $a$ must belong to exactly one set $A_{i}$ (for $i\in[L]$) since the sets $A_{i}$ are disjoint by construction. Moreover, since $N_{o}^{(i^{\prime})}\subseteq A_{i^{\prime}}$ for all $i^{\prime}\in[L]$, we further must have $N_{o_{j}}=N^{(i)}_{o_{j}}=\{a\}$, $N_{o_{j^{\prime}}}=N_{o_{j^{\prime}}}^{(i)}=\{a\}$, and $o_{j},o_{j^{\prime}}\in O^{(s)}_{i}$ for the same $i\in[L]$. Recall that the sets $\{N_{o}^{(i)}\mid o\in A_{\leq i-1}\cup(O\setminus O_{\leq i-1})\}$ were constructed to satisfy the properties of Lemma 2.2 with respect to $A_{\leq i}$ and $A_{\leq i-1}\cup(O\setminus O_{\leq i-1})$. Since $N_{o_{j^{\prime}}}^{(i)}=N_{o_{j}}^{(i)}=\{a\}$, the third such property implies that $(A_{\leq i}-a)\cup\{o_{j^{\prime}},o_{j}\}\in{\mathcal{I}}$. Additionally, since $o_{j^{\prime}}\neq o_{j}$ but $N_{o_{j^{\prime}}}^{(i)}=N_{o_{j}}^{(i)}$, the first such property implies that $o_{j^{\prime}},o_{j}\not\in A_{\leq i}\cap(A_{\leq i-1}\cup(O\setminus O_{\leq i-1}))\supseteq A_{\leq i}$. Thus, $o_{j^{\prime}},o_{j}\in O_{i}^{(s)}\setminus A_{\leq i}$ and by (1) and (2), respectively, they must obey $f(o_{j^{\prime}}\mid A_{\leq i})\geq m_{i}$ and $f(o_{j}\mid A_{\leq i}+o_{j^{\prime}})\geq m_{i}$. Thus, $(\{o_{j^{\prime}},o_{j}\},\{a\})$ is a valid $(m_{i},\varepsilon)$-improvement for $A_{\leq i}$, yielding the desired contradiction.
Claim 3.
Fix any $a\in A_{\leq L}$ and let $O^{\prime}=\{o\in O_{\leq L}\mid a\in N_{o}\}$. By the same argument as in the previous claim, since the sets $A_{i}$ are disjoint and $N_{o}^{(i)}\subseteq A_{i}$ for all $i\in[L]$, there must be some single $i\in[L]$ such that $a\in A_{i}$ and $N_{o}=N_{o}^{(i)}$ for all $o\in O^{\prime}$. For this $i$, the sets $\{N_{o}^{(i)}\mid o\in A_{\leq i-1}\cup(O\setminus O_{\leq i-1})\}$ were constructed to satisfying the properties of Lemma $2.$ By the fourth such property, $a$ appears in at most $k$ of them, and hence, $|O^{\prime}|\leq k$. ∎
4.4 Defining Auxiliary Weights Independent of $A_{\leq L}$
The weights, sets and properties given in the preceding subsections are sufficient to carry out a charging argument analogous to the one presented for the linear case in [38]. As in their proof, this charging argument would use the random shift $\tau$ to argue that, in expectation, an element $o\in O_{\leq L}\setminus O^{(s)}$ has value significantly less than all elements $a\in N_{o}$. The main difficulty in extending the argument to the submodular case is that the weights $w$ and $\bar{w}$} depend crucially on the set $A_{\leq L}$ constructed by the algorithm, and hence, also on the random choice of $\tau$. To circumvent this difficulty, we consider the incremental contribution of each element $o\in O$ to $f(O)$ alone. Formally, we order the elements $o_{1},\dotsc,o_{|O|}$ of $O$ as in the definition of $\bar{w}$}, and then, for every $o_{j}\in O$, define the auxiliary weight
$u(o_{j})\triangleq f(o_{j}\mid\{o_{1},\dotsc,o_{j-1}\})\kern 5.0pt.$
Note that $\sum_{o\in O}u(o)=f(O)-f(\emptyset)$. Moreover, the auxiliary weights $u$ are independent of the execution of Algorithm 1, and in particular, the random choice of $\tau$. The following observation relates the weights $\bar{w}$} and $u$.
Observation $4.$
For every $o_{j}\in O$, $u(o_{j})>0$ and $u(o_{j})\geq{\bar{w}}(o_{j})$. When $f$ is linear, the last inequality holds as an equality.
Proof.
Consider any $o_{j}\in O$. Since we assume that $O$ is strictly down-monotone, we must have $f(o_{j}\mid O-o_{j})>0$ for all $o_{j}\in O$. The submodularity of $f$ then implies that $0<f(o_{j}\mid O-o_{j})\leq f(o_{j}\mid\{o_{1},\dotsc,o_{j-1}\})=u(o_{j})$. Thus, $u(o_{j})>0$, and by using submodularity again, this further implies that
$u(o_{j})={}\max\{0,u(o_{j})\}$ $=\max\{0,f(o_{j}\mid\{o_{1},\dotsc,o_{j-1}\})\}$
$\geq{}\max\{0,f(o_{j}\mid(A_{\leq L}-o_{j})\cup\{o_{1},\dotsc,o_{j-1}\}\}={\bar{w}}(o_{j})\kern 5.0pt.$
Furthermore, when $f$ is linear, the inequality in the above derivation holds as an equality since both its sides are equal to $\max\{0,f(\{o_{j}\})\}$. ∎
For general submodular functions, we may have a strict inequality $u(o)>{\bar{w}}(o)$ for some $o\in O$. In the following section, this potential discrepancy results in a loss proportional to $u(o)-{\bar{w}}(o)$ in our charging analysis. In contrast, the following lemma shows that this discrepancy can also be used to derive an alternative lower bound on $f(A_{\leq L})$. To get our results, we balancing this lower bound against the bounds obtained in Section 4.5 via our charging scheme.
Lemma 4.5.
It holds that
$\sum_{o\in O}[u(o)-{\bar{w}}(o)]\leq f(A_{\leq L})-f(A_{\leq L}\mid O)\kern 5.0pt.$
Proof.
By the definitions of the weights $u(o)$ and ${\bar{w}}(o)$,
$\sum_{o\in O}[u(o)-{\bar{w}}(o)]={}$ $\sum_{j=1}^{|O|}f(o_{j}\mid\{o_{1},\dotsc,o_{j-1}\})-\sum_{o_{j}\in O}\max\{0,f(o_{j}\mid(A_{\leq L}-o_{j})\cup\{o_{1},\dotsc,o_{j-1}\})\}$
[…]
$={}$ $\left[f(O)-f(\emptyset)\right]-\left[f(O\cup A_{\leq L})-f(A_{\leq L})\right]$
$\leq{}$ $f(A_{\leq L})-f(A_{\leq L}\mid O)\kern 5.0pt.\qed$
4.5 The Charging Scheme
Recall that Section 4.3 classified the elements of $O$ into three types: the elements of $O^{(s)}$, the elements of $O\setminus O_{\leq L}$, and the remaining elements (i.e., the elements of $O_{\leq L}\setminus O^{(s)}$). Each one of the first three lemmata of this section upper bounds $u(o)$ for the elements $o\in O$ of one of these types. Since $u(o)$ is the marginal contribution of $o$ to $O$, later in the section we are able to derive a bound on $f(O)$ by combining the upper bounds on $u(o)$ proved by the three lemmata. As noted previously, all our upper bounds on $u(o)$ include a loss term that is proportional to $u(o)-{\bar{w}}(o)$. This loss term arises due to the possible discrepancy between the weights $\bar{w}$} and the auxiliary weights $u$ we are bounding. To get our final bound on $f(O)$, we need also Lemma 4.5, which gives a bound on $f(O)$ that improves as $\sum_{o\in O}[u(w)-{\bar{w}}(o)]$ grows, and thus, allows us to cancel the above-mentioned loss terms.
Lemma 4.6.
For each $o\in O^{(s)}$, $u(o)\leq(1+\varepsilon)\cdot w(N_{o})+[u(o)-{\bar{w}}(o)]$. Moreover, $\sum_{o\in O^{(s)}}w(N_{o})\leq f(A_{\leq L}\mid\emptyset)$.
Proof.
Claim 1 of Lemma 4.3 immediately implies that for any $o\in O^{(s)}$,
$u(o)={\bar{w}}(o)+[u(o)-{\bar{w}}(o)]\leq(1+\varepsilon)\cdot w(N_{o})+[u(o)-{\bar{w}}(o)]\kern 5.0pt.$
Claim 2 of Lemma 4.3 further implies that every element of $A_{\leq L}$ appears in $N_{o}$ for at most one element $o\in O^{(s)}$. Together with the fact that $w(a)>0$ for every $a\in A_{\leq L}$ (claim 1 of Lemma 4.2), this implies that
$\sum_{o\in O^{(s)}}w(N_{o})\leq\sum_{a\in A_{\leq L}}w(a)=f(A_{\leq L}\mid\emptyset)\kern 5.0pt.\qed$
Lemma 4.7.
For each $o\in O\setminus O_{\leq L}$, $u(o)=u(o)-{\bar{w}}(o)$.
Proof.
This follows immediately from claim 3 of Lemma 4.2, which states that ${\bar{w}}(o)=0$ for all $o\in O\setminus O_{\leq L}$. ∎
Next, we would like to bound $u(o)$ for the elements in $O_{\leq L}\setminus O^{(s)}$. Unlike the bounds in the previous two lemmata, this bound depends on the random shift $\tau$ selected by the algorithm. To formulate this bound, let us define, for each $o\in O$, $m^{(o)}\triangleq\min\{m_{i}\mid i\geq 0,m_{i}\geq u(o)\}$ to be the value of the smallest threshold $m_{i}$ that is at least $u(o)$ (observe that $m^{(o)}$ is well-defined since $u(o)>0$ by Observation 4.4 and $u(o)\leq W\leq m_{0}$ by submodularity). Additionally, we define the ratio $r_{o}\triangleq m^{(o)}/u(o)$. Notice that $r_{o}$ is always in the range $[1,2)$ since $u(o)\leq m^{(o)}$ and $m_{i-1}/m_{i}=2$ for all ii. Given this notation, we can now state the promised bound as Lemma 4.8. This lemma introduces a parameter $d\geq 1$, which is later used to control the balance between the bound of this lemma and the aforementioned bound from Lemma 4.5.
Lemma 4.8.
For every $o\in O_{\leq L}\setminus O^{(s)}$ and any $d\geq 1$,
$\rho_{o,d}\cdot u(o)\leq w(N_{o})+d[u(o)-{\bar{w}}(o)]\kern 5.0pt,$
where $\rho_{o,d}\triangleq r_{o}$ when $f$ is linear, and $\rho_{o,d}\triangleq\min\Big\{r_{o},\frac{1-\frac{1}{2}(1-1/d)}{1-r_{o}^{-1}(1-1/d)}\Big\}$ when $f$ is a general submodular function.
Proof.
We begin the proof by showing that for all $o\in O_{\leq L}\setminus O^{(s)}$, there exist some threshold $m_{j}$ such that ${\bar{w}}(o)\leq m_{j}\leq w(N_{o})$. Since $o\in O_{\leq L}$, we must have $o\in O_{i}$ for some unique $i\in[L]$, and thus, $N_{o}$ is well-defined and obeys $|N_{o}|\geq 1$. Since $o\not\in O^{(s)}$, this implies that either ${\bar{w}}(o)\leq m_{i}$ or $|N_{o}|\geq 2$. In the first case,
${\bar{w}}(o)\leq m_{i}\leq|N_{o}|\cdot m_{i}\leq w(N_{o})\kern 5.0pt,$
where the last inequality holds since $N_{o}\subseteq A_{i}$ and $w(a)\geq m_{i}$ for every $a\in A_{i}$ (by claim 1 of Lemma 4.2). In the second case,
${\bar{w}}(o)\leq m_{i-1}=2m_{i}\leq|N_{o}|\cdot m_{i}\leq w(N_{o})\kern 5.0pt,$
where the first inequality holds by claim 2 of Lemma 4.2 because $o\in O_{i}$, and the last inequality again holds since $N_{o}\subseteq A_{i}$ and $w(a)\geq m_{i}$ for every $a\in A_{i}$ (claim 1 of Lemma 4.2). Thus, in all cases there is indeed always a threshold $m_{j}$ with ${\bar{w}}(o)\leq m_{j}\leq w(N_{o})$.
We now consider two cases based on the relationship between $u(o)$ and the threshold $m_{j}$. If $u(o)\leq m_{j}$, then because $m^{(o)}$ is the smallest threshold that is at least $u(o)$, we get
$u(o)=\frac{m^{(o)}}{r_{o}}\leq\frac{m_{j}}{r_{o}}\leq\frac{w(N_{o})}{r_{o}}\kern 5.0pt.$
Rearranging this inequality yields
$r_{o}\cdot u(o)\leq w(N_{o})\leq w(N_{o})+d[u(o)-{\bar{w}}(o)]\kern 5.0pt,$ (3)
where the last inequality follows from $u(o)\geq{\bar{w}}(o)$ (Observation 4.4). When the function $f$ is linear, this case is the only possible case because $u(o)={\bar{w}}(o)\leq m_{j}$ for such $f$ (recall that $u(o)>0$ by Observation 4.4). This completes the proof of the lemma for linear functions.
For general submodular functions, we may also have $u(o)>m_{j}$. In this case, $m^{(o)}$ must be at least $m_{j+1}$, which implies $m_{j}=\frac{1}{2}m_{j+1}\leq\frac{1}{2}m^{(o)}=\frac{u(o)\cdot r_{o}}{2}$. Thus,
$\Biggl(1-\frac{1-1/d}{2}\Biggr)\cdot u(o)$ $=m_{j}-\frac{1-1/d}{2}\cdot u(o)+[u(o)-m_{j}]$
$\leq m_{j}-\frac{(1-1/d)}{2}\cdot\frac{2m_{j}}{r_{o}}+[u(o)-m_{j}]$
$=(1-r_{o}^{-1}(1-1/d))m_{j}+[u(o)-m_{j}]\kern 5.0pt.$
Note that, since $r_{o}\geq 1$, it must hold that $1-r_{o}^{-1}(1-1/d)\geq 1/d$. Thus, dividing both sides of the above inequality by $1-r_{o}^{-1}(1-1/d)$ and simplifying yields
$\frac{1-\frac{1}{2}(1-1/d)}{1-r_{o}^{-1}(1-1/d)}\cdot u(o)\leq m_{j}+\left(1-r_{o}^{-1}(1-1/d)\right)^{-1}[u(o)-m_{j}]\\ \leq m_{j}+d[u(o)-m_{j}]\leq w(N_{o})+d[u(o)-{\bar{w}}(o)]\kern 5.0pt,$ (4)
where we have used the fact that ${\bar{w}}(o)\leq m_{j}\leq w(N_{o})$ in the last inequality. The guarantee of the lemma for general submodular functions now follows by taking the smaller of the two lower bounds on $w(N_{o})+d[u(o)-{\bar{w}}(o)]$ given by (3) and (4) for the two cases. ∎
Combining all the above bounds, we obtain the following proposition.
Proposition 4.9.
Assuming $f$ is a submodular function such that $f(\emptyset)\geq 0$, the output set $A_{\leq L}$ of Algorithm 1 obeys, for every $d\geq d^{\prime}\geq 1$ and any strictly down-monotone $O\in{\mathcal{I}}$,
[…]
where $\rho_{o,d}\triangleq r_{o}$ when $f$ is linear and $\rho_{o,d}\triangleq\min\Big\{r_{o},\frac{1-\frac{1}{2}(1-1/d)}{1-r_{o}^{-1}(1-1/d)}\Big\}$ when $f$ is a general submodular function.
Proof.
Combining $d^{\prime}$ times the bounds from Lemmata 4.6 and 4.7 with the bound from Lemma 4.8, we obtain
[…]
$\leq\sum_{\mathclap{o\in O_{\leq L}}}w(N_{o})+((1+\varepsilon)d^{\prime}-1)\sum_{\mathclap{o\in O^{(s)}}}w(N_{o})+d\cdot\sum_{o\in O}[u(o)-{\bar{w}}(o)]$
$\leq\sum_{\mathclap{a\in A_{\leq L}}}k\cdot w(a)+((1+\varepsilon)d^{\prime}-1)\sum_{\mathclap{o\in O^{(s)}}}w(N_{o})+d\cdot\sum_{o\in O}[u(o)-{\bar{w}}(o)]$
$\leq(k+(1+\varepsilon)d^{\prime}-1)\cdot f(A_{\leq L}\mid\emptyset)+d\cdot\sum_{o\in O}[u(o)-{\bar{w}}(o)]\kern 5.0pt,$
where the second inequality follows from the condition $d\geq d^{\prime}\geq 1$ and the inequality $u(o)-{\bar{w}}(o)\geq 0$ (which holds by Observation 4.4); the third inequality follows from the fact that each $a\in A_{\leq L}$ appears in at most $k$ sets $N_{o}$ (claim 3 of Lemma 4.3) and obeys $w(a)\geq 0$ (claim 1 of Lemma 4.2); and the final inequality follows from the inequality $\sum_{o\in O^{(s)}}w(N_{o})\leq f(A_{\leq L}\mid\emptyset)$ proved by Lemma 4.6.
Taking expectations on both sides, and noting that $f(O)$ and $u(o)$ do not depend on the randomness of the algorithm, we obtain
[…]
$\leq d^{\prime}\cdot f(\emptyset)+(k+(1+\varepsilon)d^{\prime}-1)\cdot{\mathbb{E}}[f(A_{\leq L}\mid\emptyset)]$
[…]
Rearranging this inequality gives the claimed result. ∎
Using Proposition 4.9 we can now prove our main results for linear and monotone submodular functions. We begin with the simpler case of linear functions.
Theorem 4.10.
When $f$ is a linear function, the approximation guarantee of Algorithm 1 is at most $k\ln 2+1-\ln 2+\varepsilon$.
Proof.
For linear functions, $u(o)={\bar{w}}(o)=f(\{o\})$ for all $o\in O$, and hence, the terms ${\mathbb{E}}[u(o)-{\bar{w}}(o)]$ in Proposition 4.9 all vanish. Then, since $\rho_{o,d}=r_{o}$ for linear functions, we have
[…]
Let $d^{\prime}=\ln^{-1}2$. Below, we prove that ${\mathbb{E}}[r_{o}]=d^{\prime}$ for every $o\in O$, which implies that the last term on the rightmost side of the above inequality is 0. Rearranging the above inequality then yields
${\mathbb{E}}[f(A_{\leq L})]\geq\frac{d^{\prime}}{k+(1+\varepsilon)d^{\prime}-1}\cdot f(O)=\frac{\ln^{-1}2}{k+(1+\varepsilon)\ln^{-1}2-1}\cdot f(O)=\frac{f(O)}{k\ln 2+1-\ln 2+\varepsilon}\kern 5.0pt.$
Recalling that there must exist at least one optimal solution $O$ that is also strictly down-monotone, the theorem follows from this inequality. The rest of the proof is devoted to proving that indeed ${\mathbb{E}}[r_{o}]=d^{\prime}$.
Recall that for each $o\in O$, $r_{o}$ is the ratio $m^{(o)}/u(o)$, where $m^{(o)}=\min\{m_{i}\mid i\geq 0,m_{i}\geq u(o)\}$ is the lowest value of a threshold $m_{i}$ that is at least $u(o)$. We also recall that, for each $i$, $m_{i}=W\tau 2^{-i}=W2^{-i+\alpha}$, where $\alpha$ is chosen uniformly at random from $(0,1]$. Intuitively, this implies that $r_{o}=2^{\beta}$ for some $\beta\in[0,1)$ that is obtained by cyclically shifting $\alpha$, and is thus, distributed uniformly at random from $[0,1)$. We give a formal proof of this fact in Lemma B.1 in Appendix B. Given this fact,
${\mathbb{E}}\left[r_{o}\right]=\int_{0}^{1}2^{\beta}\,\mathrm{d}\beta=\frac{2^{\beta}}{\ln 2}\biggr|_{\beta=0}^{\beta=1}=\ln^{-1}2=d^{\prime}\kern 5.0pt.\qed$
The following proposition gives a guarantee for Algorithm 1 when the objective is a general submodular (rather than linear) function. Our main result for monotone submodular functions readily follows from this proposition (see Corollary 4.12), but the proposition itself applies also to non-monotone submodular functions, and is used in Setion 5 to derive our main result for such functions.
Proposition 4.11.
When $f$ is a non-negative submodular function, the output set $A_{\leq L}$ of Algorithm 1 obeys, for every $d\geq 1$ and any strictly down-monotone $O\in{\mathcal{I}}$,
$(k+(1+\varepsilon)d^{\prime}+d-1)\cdot{\mathbb{E}}[f(A_{\leq L})]\geq d^{\prime}\cdot f(O)+d\cdot{\mathbb{E}}[f(A_{\leq L}\mid O)]\kern 5.0pt,$
where $d^{\prime}=\frac{1-1/d}{2\ln 2}+\frac{d+1}{2d}$.
Proof.
One can verify that $d\geq d^{\prime}\geq 1$ for every $d\geq 1$. Proposition 4.9 and Lemma 4.5 thus give together
[…]
Below we prove that ${\mathbb{E}}[\rho_{o,d}]=d^{\prime}$ for every $o\in O$, which implies that the last term on the rightmost side of the above inequality is 0. Notice that, given this observation, the proposition follows by using the linearity of expectation and rearranging the resulting inequality. The rest of the proof is devoted to proving that indeed ${\mathbb{E}}[\rho_{o,d}]=d^{\prime}$.
Recall that Lemma B.1 in Appendix B proves that $r_{o}=2^{\beta}$ for $\beta$ that is distributed uniformly at random from $[0,1)$. This property of $r_{o}$ underlies the calculation below.
${\mathbb{E}}[\rho_{o,d}]$ $={\mathbb{E}}\left[\min\left\{r_{o},\frac{1-\frac{1}{2}(1-1/d)}{1-r_{o}^{-1}(1-1/d)}\right\}\right]$
$=\int_{0}^{1}\min\left\{2^{\beta},\frac{1-\frac{1}{2}(1-1/d)}{1-2^{-\beta}(1-1/d)}\right\}\,\mathrm{d}\beta$
$=\int_{0}^{\log_{2}(1+\frac{1}{2}(1-1/d))}2^{\beta}\,\mathrm{d}\beta+\int_{\log_{2}(1+\frac{1}{2}(1-1/d))}^{1}\frac{1-\frac{1}{2}(1-1/d)}{1-2^{-\beta}(1-1/d)}\mathrm{d}\beta$
$=\frac{2^{\beta}}{\ln 2}\biggr|_{\beta=0}^{\beta=\log_{2}(1+\frac{1}{2}(1-1/d))}+\frac{1-\frac{1}{2}(1-1/d)}{\ln 2}\cdot\ln\left(2^{\beta}-(1-1/d)\right)\biggr|_{\beta=\log_{2}(1+\frac{1}{2}(1-1/d))}^{\beta=1}$
$=\frac{1-1/d}{2\ln 2}+\frac{1-\frac{1}{2}(1-1/d)}{\ln 2}\cdot\left(\ln(1+1/d)-\ln(\tfrac{1}{2}(1+1/d))\right)$
$=\frac{1-1/d}{2\ln 2}+1-\tfrac{1}{2}(1-1/d)=\frac{1-1/d}{2\ln 2}+\frac{d+1}{2d}=d^{\prime}\kern 5.0pt.\qed$
Corollary 4.12.
When $f$ is a non-negative monotone submodular function, the approximation guarantee of Algorithm 1 is at most $\frac{2k\ln 2}{1+\ln 2}+O(\sqrt{k})\leq 0.819k+O(\sqrt{k}).$
Proof.
In the case that $f$ is monotone, the expression $f(A_{\leq L}\mid O)$ in Proposition 4.11 is always non-negative. Thus, for any $d\geq 1$, we have
$(k+(1+\varepsilon)d^{\prime}+d-1)\cdot{\mathbb{E}}[f(A_{\leq L})]\geq d^{\prime}\cdot f(O)\kern 5.0pt,$
where $d^{\prime}=\frac{1-1/d}{2\ln 2}+\frac{d+1}{2d}$. Substituting $d^{\prime}$ into this inequality, we get that the ratio $f(O)/{\mathbb{E}}[f(A_{\leq L})]$ is at most
$\frac{k+d+(1+\varepsilon)\cdot\Big(\frac{1-1/d}{2\ln 2}+\frac{d+1}{2d}\Big)-1}{\frac{1-1/d}{2\ln 2}+\frac{d+1}{2d}}=1+\varepsilon+\frac{k+d-1}{\frac{1-1/d}{2\ln 2}+\frac{d+1}{2d}}\\ =1+\varepsilon+\frac{k+d-1}{\frac{1+\ln 2}{2\ln 2}-d^{-1}\cdot\frac{1-\ln 2}{2\ln 2}}\leq 1+\varepsilon+\frac{(k+d-1)\cdot\big(1+2d^{-1}\cdot\frac{1-\ln 2}{1+\ln 2}\big)}{\frac{1+\ln 2}{2\ln 2}}\kern 5.0pt,$
where the inequality uses the fact that for every two positive values $a,b$ obeying $2b\leq a$, we have $\frac{1}{a-b}\leq\frac{1+2b/a}{a}$. Selecting $d=\sqrt{k}$, and recalling that $\varepsilon<1$, makes the above expression equal to $\frac{2k\ln 2}{1+\ln 2}+O(\sqrt{k})$, which proves the corollary since there must exist some optimal solution $O$ that is also strictly down-monotone. ∎
[…]
5 Algorithm for Non-monotone Submodular Functions
In this section, we present and analyze Algorithm 2, which we use to prove Theorem $1.$ This algorithm receives as input a non-negative submodular function $f\colon 2^{E}\to{{\mathbb{R}}_{\geq 0}}$, a matroid $k$-parity constraint $(E,{\mathcal{I}})$, and a positive integer parameter $\ell$. Algorithm 2 runs for $\ell$ iterations. In each iteration $i$, Line 2 considers the restriction of our problem to some ground set $E_{i-1}\subseteq E$. Note that the restricted $f_{i}$ inherits the submodularity of $f$ and the resulting constraint $(E_{i-1},2^{E_{i-1}}\cap{\mathcal{I}})$ is also a matroid $k$-parity constraint. In each iteration, we apply Algrithm 1 to obtain a feasible set $B_{i}\subseteq E_{i-1}$. Next, in Line 2 the algorithm invokes the Double Greedy algorithm of Buchbinder et al. [5]. This algorithm takes a non-negative submodular function $f\colon 2^{E}\to{{\mathbb{R}}_{\geq 0}}$ and a set $S\subseteq E$, and returns a set $T\subseteq S$ such that ${\mathbb{E}}[f(T)]\geq\frac{1}{2}\cdot\max_{S^{\prime}\subseteq S}f(S^{\prime})$. We denote by $B^{\prime}_{i}$ the output of Double Greedy applied to our set $B_{i}$. Algorithm 2 then returns as its output the best set constructed amongst all the sets $B_{i}$ and $B^{\prime}_{i}$ in all $\ell$ iterations.
1Let $E_{0}\leftarrow E$.
2 for $i=1$ to $\ell$ do
3   Execute Algorithm 1 on the function $f$, the constraint $(E_{i-1},2^{E_{i-1}}\cap{\mathcal{I}})$, and an arbitrary constant value $\varepsilon\in(0,1)$. Let $B_{i}$ be the output set of this algorithm.
4   Execute the Double Greedy algorithm of [5] on the function $f$ and the set $B_{i}$, and let $B^{\prime}_{i}$ be the output set of this algorithm.
5   Let $E_{i}\leftarrow E_{i-1}\setminus B_{i}$.
return the set $B^{*}\in\{B_{i}\mid i\in[\ell]\}\cup\{B^{\prime}_{i}\mid i\in[\ell]\}$ with maximum value $f(B^{*})$.
Algorithm 2 Repetitions Algorithm $(E,f,{\mathcal{I}},\ell)$
We begin the analysis of Algorithm 2 with the following two lemmata, which lower bound the values of the individual sets $B_{i}$ and $B^{\prime}_{i}$. As in our analysis of Algorithm 1 in Section 4, we assume that $O$ is an arbitrary strictly down-monotone feasible solution, and let $B^{*}$ be the set produced by Algorithm 2.
Lemma 5.1.
For every $i\in[\ell]$ and $d\geq 1$,
$(k+d+(1+\varepsilon)d^{\prime}-1)\cdot{\mathbb{E}}[f(B_{i})]\geq{}$ $d^{\prime}\cdot{\mathbb{E}}[f(O\cap E_{i-1})]+d\cdot\operatorname*{\mathbb{E}}[f(B_{i}\mid O)]\kern 5.0pt,$
where $d^{\prime}=\frac{1-1/d}{2\ln 2}+\frac{d+1}{2d}$.
Proof.
Notice that the set $B_{i}$ is the output set of the execution of Algorithm 1 invoked by Line 2 of Algorithm 2. The ground set of the matroid $k$-parity constraint passed to this execution is $E_{i-1}$, and thus, $B_{i}\subseteq E_{i-1}$. Conditioned on $E_{i-1}$, Proposition 4.11 implies that
$(k+d+(1+\varepsilon)d^{\prime}-1)\cdot{\mathbb{E}}[f(B_{i})]\geq{}$ $d^{\prime}\cdot f(O^{\prime})+d\cdot\operatorname*{\mathbb{E}}[f(B_{i}\mid O^{\prime})]$ (5)
$\geq{}$ $d^{\prime}\cdot f(O^{\prime})+d\cdot\operatorname*{\mathbb{E}}[f(B_{i}\mid O^{\prime}\cup(O\setminus E_{i-1}))]$
for every set $O^{\prime}$ that is a feasible and strictly down-monotone subset of $E_{i-1}$. The second inequality in Inequality (5) follows from the submodularity of $f$ because $B_{i}\subseteq E_{i-1}$. Now, consider the set $O\cap E_{i-1}\subseteq O$. For all $x\in O\cap E_{i-1}$, we have $f(x\mid O\cap E_{i-1}-x)\geq f(x\mid O-x)>0$ due to the submodularity of $f$ and the strict down-monotonicity of $O$. Thus, $O\cap E_{i-1}$ must also be a strictly down-monotone solution. Letting $O^{\prime}=O\cap E_{i-1}$ in (5), and then taking the expectation over the set $E_{i-1}$, completes the proof of our claim. ∎
Lemma 5.2.
For every $i\in[\ell]$, ${\mathbb{E}}[f(B^{\prime}_{i})]\geq\frac{1}{2}\cdot{\mathbb{E}}[f(O\cap B_{i})]$.
Proof.
By the properties of the Double Greedy algorithm, conditioned on the set $B_{i}$, ${\mathbb{E}}[f(B^{\prime}_{i})]\geq\frac{1}{2}\cdot f(O\cap B_{i})$. The lemma follows by taking the expectation of both sides of this inequality over the set $B_{i}$. ∎
Combining the above results, we have the following. Recall that $B^{*}$ denotes the output set of Algorithm 2.
Corollary 5.3.
For every $i\in[\ell]$ and $d\geq 1$,
$(k+d+(2i-1+\varepsilon)d^{\prime}-1)\cdot{\mathbb{E}}[f(B^{*})]\geq d^{\prime}\cdot{\mathbb{E}}[f(O)]+d\cdot\operatorname*{\mathbb{E}}[f(B_{i}\mid O)]\kern 5.0pt,$
where $d^{\prime}=\frac{1-1/d}{2\ln 2}+\frac{d+1}{2d}$.
Proof.
By the definition of $B^{*}$,
$(k+d+(1+\varepsilon)d^{\prime}-1)\cdot{\mathbb{E}}[f(B^{*})]\geq{}$ $(k+d+(1+\varepsilon)d^{\prime}-1)\cdot{\mathbb{E}}[f(B_{i})]$
$\geq{}$ $d^{\prime}\cdot{\mathbb{E}}[f(O\cap E_{i-1})]+d\cdot\operatorname*{\mathbb{E}}[f(B_{i}\mid O)]$
$\geq{}$ $d^{\prime}\cdot{\mathbb{E}}\bigg[f(O)-\sum_{j=1}^{i-1}f(O\cap B_{j})\bigg]+d\cdot\operatorname*{\mathbb{E}}[f(B_{i}\mid O)]$
$={}$ $d^{\prime}\cdot f(O)-d^{\prime}\cdot\sum_{j=1}^{i-1}{\mathbb{E}}[f(O\cap B_{j})]+d\cdot\operatorname*{\mathbb{E}}[f(B_{i}\mid O)]$
$\geq{}$ $d^{\prime}\cdot f(O)-2d^{\prime}\cdot\sum_{j=1}^{i-1}{\mathbb{E}}[f(B^{\prime}_{j})]+d\cdot\operatorname*{\mathbb{E}}[f(B_{i}\mid O)]$
$\geq{}$ $d^{\prime}\cdot f(O)-2d^{\prime}(i-1)\cdot{\mathbb{E}}[f(B^{*})]+d\cdot\operatorname*{\mathbb{E}}[f(B_{i}\mid O)]\kern 5.0pt,$
where the second inequality holds by Lemma 5.1, the third inequality follows from the submodularity of $f$ since $O\cap B_{1},O\cap B_{2},\dotsc,O\cap B_{i-1}$ and $O\cap E_{i-1}$ are a disjoint partition of $O$, and the penultimate inequality holds by Lemma 5.2. The corollary now follows by rearranging this inequality. ∎
[…]
Proof.
Summing the guarantee of Corollary 5.3 over all $i\in[\ell]$, we get
$\ell\cdot(k+d+(\ell+\varepsilon)d^{\prime}-1)\cdot{\mathbb{E}}[f(B^{*})]={}$ $\sum_{i=1}^{\ell}(k+d+(2i-1+\varepsilon)d^{\prime}-1)\cdot{\mathbb{E}}[f(B^{*})]$
$\geq{}$ $d^{\prime}\ell\cdot f(O)+d\cdot\operatorname*{\mathbb{E}}\bigg[\sum_{i=1}^{\ell}f(B_{i}\mid O)\bigg]$
$\geq{}$ $d^{\prime}\ell\cdot f(O)+d\cdot\operatorname*{\mathbb{E}}[f((\cup_{i=1}^{\ell}B_{i})\mid O)\bigg]$
$\geq{}$ $(d^{\prime}\ell-d)\cdot f(O)\kern 5.0pt,$
where the second inequality follows from the submodularity of $f$, and the last inequality follows from $f$’s non-negativity. Recall that there must be some strictly down-montone set $O$ that is also an optimal solution. Thus, the approximation ratio of Algorithm 2 is at most
$\frac{\ell\cdot(k+d+(\ell+\varepsilon)d^{\prime}-1)}{d^{\prime}\ell-d}={}$ $\ell+\varepsilon+\frac{\ell\cdot(k+2d-1)+\varepsilon d}{d^{\prime}\ell-d}$
$={}$ $\ell+\varepsilon+\frac{k+2d-1+\varepsilon d/\ell}{\frac{1-1/d}{2\ln 2}+\frac{d+1}{2d}-d/\ell}=\ell+\varepsilon+\frac{k+2d-1+\varepsilon d/\ell}{\frac{1+\ln 2}{2\ln 2}-d^{-1}\cdot\frac{1-\ln 2}{2\ln 2}-d/\ell}$
$\leq{}$ $\ell+\varepsilon+\frac{(k+2d-1+\varepsilon d/\ell)\cdot\big(1+2d^{-1}\cdot\frac{1-\ln 2}{1+\ln 2}+\frac{2d}{\ell}\cdot\frac{2\ln 2}{1+\ln 2}\big)}{\frac{1+\ln 2}{2\ln 2}}\kern 5.0pt,$
where the inequality holds for $\ell\geq 3d$ because for every two positive values $a,b$ obeying $2b\leq a$, it holds that $\frac{1}{a-b}\leq\frac{1+2b/a}{a}$. Choosing now $\ell=\lceil 3k^{2/3}\rceil$ and $d=k^{1/3}$, and recalling that $\varepsilon$ is assigned a constant value from $(0,1)$, we get that the approximation ratio of Algorithm 2 is at most $\frac{2k\ln 2}{1+\ln 2}+O(k^{2/3})$.
It remains to bound the time complexity of Algorithm 2. The Double Greedy algorithm of [5] runs in $O(|E|)$ time, and in Appendix A, we show that it is possible to simulate Algorithm 1 in $O(\varepsilon^{-1}|E|^{4})$ time. Since Algorithm 2 invokes each one of these algorithms $\ell=\lceil 3k^{2/3}\rceil$ times, we get that Algorithm 2 can be implemented to run in
$O(k^{2/3}\varepsilon^{-1}|E|^{4})=O(k^{2/3}|E|^{4})=\operatorname{Poly}(|E|,k)$
time. When $k\leq 2|E|$, this implies the theorem. When $k>2|E|$, the theorem still holds since in this case the naïve algorithm that simply returns the best feasible singleton from $E$ already obeys all the properties guaranteed by the theorem. ∎
Appendix A Polynomial Time Implementation of Algorithm 1
In Section 3, we proved only that Algorithm 1 terminated. Here, we show that it is possible to efficiently implement Algorithm 1 in polynomial time. Recall that $L$ is the number of iterations that Algorithm 1 performs. In Section 3, we have only shown (in Observation 3.1) that $L$ is finite. One can note, however, that the number of iterations $i$ constructing a non-empty set $A_{i}$ is limited to $|E|$ because the sets $A_{i}$ are disjoint. Moreover, immediately after each iteration that constructs a non-empty set $A_{i^{\prime}}$, and before starting any further iteration, we can already calculate the index $i$ of the next iteration in which $A_{i}$ will be non-empty. Thus, it is possible to accelerate the algorithm by skipping directly to iteration $i$, rather then going through all the iterations $i^{\prime}+1,i^{\prime}+2,\dotsc,i-1$, and generating an empty set in each one of them. Formally, the index $i$ to which we should skip after iteration $i^{\prime}$ is given by
$i=\lceil\log_{2}(W\tau)-\log_{2}w\rceil\kern 5.0pt,$
where $w$ is the maximum marginal contribution $f(e\mid A_{\leq i^{\prime}})$ of any element $e$ with $A_{\leq i^{\prime}}+e\in{\mathcal{I}}$. This formula is well-defined only for $w>0$, but this is not a problem since the termination condition of Algorithm 1 is equivalent to terminating when $w\leq 0$. The accelerated version of Algorithm 1 obtained by implementing this observation is shown in Algorithm 3.
1Let $W\leftarrow\max_{e\in E}f(\{e\}\mid\emptyset)$.
2 Let $\alpha$ be a uniformly random value from $(0,1]$, and let $\tau\leftarrow 2^{\alpha}$.
3 Define $m_{i}\triangleq W\tau 2^{-i}$ for every integer $i\geq 0$.
4
51exLet $i\leftarrow 0$ and $A_{\leq 0}\leftarrow\emptyset$.
6 while there exists $e\in E\setminus A_{\leq i}$ such that $f(e\mid A_{\leq i})>0$ and $A\leq i+e\in{\mathcal{I}}$ do
   Let $i^{\prime}\leftarrow i$
 $\triangleright$ Store the value of $i$ from the previous iteration.
7   Let $E_{i^{\prime}}\leftarrow\{e\in E\setminus A_{\leq i^{\prime}}\mid A_{\leq i^{\prime}}+e\in{\mathcal{I}}\}$.
8   Let $w_{i^{\prime}}\leftarrow\max_{e\in E_{i^{\prime}}}f(e\mid A_{\leq i^{\prime}})$.
9   Update $i\leftarrow\lceil\log_{2}(W\tau)-\log_{2}w_{i^{\prime}}\rceil$.
10   Let $A_{i}\leftarrow\emptyset$.
11 while there is some $(m_{i},\varepsilon)$-improvement $(S,N)$ for $A_{\leq i^{\prime}}\cup A_{i}$ with $N\subseteq A_{i}$ do
12    $A_{i}\leftarrow(A_{i}\setminus N)\cup S$.
13
14  Let $A_{\leq i}\leftarrow A_{\leq i^{\prime}}\cup A_{i}$.
15
16return $A_{\leq i}$
Algorithm 3 Efficient Greedy/Local-Search Hybrid Algorithm $(E,f,{\mathcal{I}},\varepsilon)$
Similarly to Algorithm 1, the condition on Line 3 of Algorithm 3 implies that the set $A_{\leq i}=A_{\leq i^{\prime}}\cup A_{i}$ finalized at the end of each iteration carried out by the algorithm must be locally optimal with respect to $(m_{i},\varepsilon)$-improvements. In particular, the following direct analogue of Observation 3.2 holds.
Observation A.1.
For every iteration $i$ in which Algorithm 3 produces a set $A_{i}$, we have $f(e\mid A_{\leq i})<m_{i}$ for all $e\in E\setminus A_{\leq i}$ that obey $A_{\leq i}+e\in{\mathcal{I}}$ at the end of iteration ii.
Our next objective is to show that Algorithm 3 correctly simulates Algorithm 1. We assume in the proof that the two algorithms use the same method for finding the next $(m_{i},\varepsilon)$-improvement in each iteration of the inner loop, and we fix the values of any random bits used by this method (if it is randomized), as well as the value of $\alpha$. Once these values are fixed, showing that Algorithm 3 correctly simulates Algorithm 1 boils down to proving that they both return the same output set. Intuitively, to prove this, we need to show that the “fast forward” operation done by Line 3 of Algorithm 3 works as expected by the intuition given at the beginning of this appendix. The next lemma is the first step towards this goal. It proves the simpler claim that the fast forward operation always increases ii. Notice that this lemma is necessary for guaranteeing that each iteration of Algorithm 3 has a unique index ii.
Lemma A.2.
Every execution of Line 3 of Algorithm 3 strictly increases the value of ii. Moreover, after $i$ is increased, we have $m_{i}\leq w_{i^{\prime}}<m_{i-1}=2m_{i}$
Proof.
In the first execution of Line 3, $w_{0}=W$, which implies that $i$ is increased from 0 to
$\lceil\log_{2}(W\tau)-\log_{2}W\rceil=\lceil\log_{2}\tau\rceil=\lceil\alpha\rceil=1\kern 5.0pt.$
Consider now an arbitrary execution of Line 3 after the first. At the start of the iteration of this execution, Algorithm 3 considers the set $E_{i^{\prime}}=\{e\in E\setminus A_{\leq i^{\prime}}\mid A_{\leq i^{\prime}}+e\in{\mathcal{I}}\}$, where $A_{\leq{i^{\prime}}}$ is the set produced by the previous iteration. By Observation A.1, we must have $f(e\mid A_{\leq i^{\prime}})<m_{i^{\prime}}$ for all $e\in E_{i^{\prime}}$, and hence, $w_{i^{\prime}}<m_{i^{\prime}}$. Thus,
$\log_{2}(W\tau)-\log_{2}w_{i^{\prime}}>\log_{2}(W\tau)-\log_{2}m_{i^{\prime}}=\log_{2}(W\tau)-\log_{2}(W\tau 2^{-i^{\prime}})=i^{\prime}\kern 5.0pt.$
The new value assigned to $i$ by Line 3 is the ceiling of the term on the left, which must be at least $i^{\prime}+1$ since $i^{\prime}$ is an integer.
For the second part of the claim, note that by definition of $m_{i}$,
$m_{i}=W\tau 2^{-i}=W\tau 2^{-\lceil\log_{2}(W\tau)-\log_{2}w_{i^{\prime}}\rceil}\leq W\tau 2^{\log_{2}w_{i^{\prime}}-\log_{2}(W\tau)}=w_{i^{\prime}}\kern 5.0pt.$
Similarly,
$m_{i-1}=W\tau 2^{-(i-1)}=W\tau 2^{-(\lceil\log_{2}(W\tau)-\log_{2}w_{i^{\prime}}\rceil-1)}>W\tau 2^{\log_{2}w_{i^{\prime}}-\log_{2}(W\tau)}=w_{i^{\prime}}\kern 5.0pt,$
where the strict inequality holds because Algorithm 3 terminates without making a single iteration when $W\leq 0$. ∎
In Section 3, we argue that each iteration of Algorithm 1 must terminate. This argument applies also to Algorithm 3. Let $C$ be the set of all values from $\ell$] that are assigned to $i$ by Algorithm 3 at some point. Clearly, Algorithm 3 constructs sets $A_{j}$ and $A_{\leq j}$ for each $j\in C$, while Algorithm 1 constructs sets $A_{j}$ and $A_{\leq j}$ for every $j\in[L]$. The following lemma gives additional properties of the set $C$.
Lemma A.3.
For all $j\in[\ell]$, $j\in C$ if and only if Algorithm 1 produces a set $A_{j}\neq\emptyset$. Moreover, for all $j\in C\cup\{0\}$, Algorithms 2 and 3 construct the same set $A_{\leq j}$.
Proof.
We proceed by induction on $j$. For $j=0$, the lemma simply states that $A_{\leq 0}$ is identical under both Algorithm 1 and Algorithm 3, which is the case since both algorithms set it to be the empty set.
Suppose $j\geq 1$, and let $j^{\prime}<j$ be the next largest index in $C\cup\{0\}$. By the induction hypothesis, both algorithms agree on the value of $A_{\leq j^{\prime}}$, and for any iteration $i$ with $j^{\prime}<i<j$, Algorithm 1 must have $A_{i}=\emptyset$. Thus, $A_{\leq j-1}=A_{\leq j^{\prime}}\cup(\cup_{i=j^{\prime}+1}^{j-1}A_{i})=A_{\leq j^{\prime}}$. We now need to consider two cases, depending on whether $j\in C$ or not. Assume first that $j\in C$. Then, both algorithms produce the same set $A_{j}$ because (i) $A_{\leq j-1}=A_{\leq j^{\prime}}$, (ii) we fixed the value of $\alpha$, and (iii) we assume that both algorithms use the same method for choosing $(m_{i},\varepsilon)$-improvements and we fixed the random bits used by this method (if any). This implies that both algorithms set $A_{\leq j}$ to the same value because Algorithm 1 sets $A_{\leq j}$ to $A_{j}\cup A_{\leq j-1}$ and Algorithm 3 sets $A_{\leq j}$ to $A_{j}\cup A_{\leq j^{\prime}}$. We now show that $A_{j}\neq\emptyset$. Since $j\in C$, Algorithm 3 has $w_{j^{\prime}}\geq m_{j}$ by Lemma A.2. Thus, there must be some $x\in E\setminus A_{\leq j^{\prime}}$ with $A_{\leq j^{\prime}}+x\in{\mathcal{I}}$ and $f(x\mid A_{\leq j^{\prime}})\geq m_{i}$. If $A_{j}=\emptyset$, then $A_{\leq j}=A_{\leq j^{\prime}}\cup\emptyset=A_{\leq j^{\prime}}$, and hence, it also holds that $A_{\leq j}+x\in{\mathcal{I}}$ and $f(x\mid A_{\leq j})\geq m_{j}$, contradicting local optimality at the end of the iteration (Observation A.1).
Consider now the case that $j\not\in C$, and assume towards a contradiction that Algorithm 1 produces a set $A_{j}\neq\emptyset$. This means that Algorithm 1 performed at least one $(m_{j},\varepsilon)$-improvement in iteration $j$, and thus, there exists an element $x\in E\setminus A_{\leq j-1}=E\setminus A_{\leq j^{\prime}}$ such that $A_{\leq j^{\prime}}+x=A_{\leq j-1}+x\in{\mathcal{I}}$ and $f(x\mid A_{\leq j^{\prime}})=f(x\mid A_{\leq j-1})\geq m_{j}$. Let us denote now by $\hat{\jmath}$ the value that $i$ gets in Algorithm 3 immediately after the value $j^{\prime}$. The existence of $x$ guarantees that such a value $\hat{\jmath}$ exists, and by the definition of $j^{\prime}$, $\hat{\jmath}$ must be at least $j+1$. Thus, by Lemma A.2,
$w_{j^{\prime}}<m_{\hat{j}-1}\leq m_{j}\leq f(x\mid A_{\leq j^{\prime}})\kern 5.0pt,$
which is a contradiction since the definition $w_{j^{\prime}}$ guarantees that $w_{j^{\prime}}\geq f(x\mid A_{\leq j^{\prime}})$. ∎
Lemma A.4.
Algorithm 3 returns the same set $A_{\leq L}$ as Algorithm 1, and the final value of $i$ in this algorithm is $L$.
Proof.
We begin by proving that both algorithms produce the same set $A_{\leq L}$. If $L=0$, then by definition, both algorithms set $A_{\leq L}=A_{\leq 0}=\emptyset$. Otherwise, if $L\geq 1$, then by Observation 3.1, Algorithm 1 constructs a non-empty set $A_{L}$, and thus, by Lemma A.3, Algorithm 3 constructs the same set $A_{\leq L}$ as Algorithm 1.
Let us now explain why the lemma follows from the fact that both algorithms produce the same set $A_{\leq L}$. Observation 3.1 shows that $f(e\mid A_{\leq L})\leq 0$ for all $e\in E\setminus A_{\leq L}$ with $A_{\leq L}+e\in{\mathcal{I}}$. Thus, the main loop of Algorithm 3 must terminate immediately after the iteration constructing $A_{\leq L}$ (or before the first iteration if $L=0$), which makes this algorithm return $A_{\leq L}$. ∎
We now get to our final result regarding Algorithm 3
Proposition A.5.
Algorithm 3 has the same output distribution as Algorithm 1, and it runs in $O(\varepsilon^{-1}|E|^{4})$ time.
Proof.
The first part of the proposition holds since Lemma A.4 guarantees that the two algorithms produce the same output set when they draw the same random bits. Thus, in the rest of this proof, we concentrate on proving that Algorithm 1 can be implemented to run in $O(\varepsilon^{-1}|E|^{4})$ time.
By Lemma A.4, $i$ never reaches any value larger than $L$. Together with the definition of $C$, this implies that the number of iterations Algorithm 3 performs is $|C|$. By Lemma A.3, each value $i\in C$ corresponds to a non-empty set $A_{i}$ in Algorithm 1, and since these sets are disjoint by construction, $|C|\leq|E|$. Thus, Algorithm 3 makes at most $|E|$ iterations. Each such iteration takes at most $O(|E|)$ time (the time required to compute $w_{i}$ and update ii) plus the time required to search and apply the $(m_{i},\varepsilon)$ improvements. Thus, the running time of Algorithm 3 is $O(|E|^{2})$ plus the total time spent searching for and applying these improvements across all iterations.
Let us denote by NN the number of improvements found throughout the algorithm. Notice that in each iteration of Algorithm 3, the number of times that the algorithms looks for an improvement is larger by exactly 11 compared to the number of improvements found in this iteration. Thus, Algorithm 3 looks for improvements at most $M+|E|$ times. Looking for an improvement can be done in $O(|E|^{3})$ time by simply enumerating over all the possible options for a set $S\subseteq E\setminus(A_{i}\cup A_{\leq i^{\prime}})$ of size at most 22 and a set $N\subseteq A_{i}$ of size at most 11. Once an improvement is found, applying it (i.e., replacing $A_{i}$ with $(A_{i}\setminus N)\cup S$) can be done in $O(1)$ time, and thus, by the above discussion, the time complexity of Algorithm 3 is
[…]
It remains to bound $M$. Consider the results of applying each of the three types of improvements in Definition 3.1 in some iteration ii.
  • Applying the first type of improvement increases $|A_{i}|$ by 1 and increases $f(A_{\leq i^{\prime}}\cup A_{i})$ by at least $\theta>0$.
  • Applying the second type of improvement does not change $|A_{i}|$ and increases $f(A_{\leq i^{\prime}})$ by at least $\varepsilon\theta$.
  • Applying the third type of improvement increases $|A_{i}|$ by 1 and does not decrease the value of $A_{\leq i^{\prime}}\cup A_{i}$. This latter claim follows from the submodularity of $f$ and Lemma A.2, which together give
    $f((A_{\leq i^{\prime}}\cup A_{i}\cup S)\setminus N)-f(A_{\leq i^{\prime}}\cup A_{i})\\ \begin{aligned} &=f(x_{1}\mid A_{\leq i^{\prime}}\cup A_{i}-y)+f(x_{2}\mid A_{\leq i^{\prime}}\cup A_{i}-y+x_{1})-f(y\mid A_{\leq i^{\prime}}\cup A_{i}-y)\\ &\geq f(x_{1}\mid A_{\leq i^{\prime}}\cup A_{i})+f(x_{2}\mid A_{\leq i^{\prime}}\cup A_{i}+x_{1})-f(y\mid A_{\leq i^{\prime}})\\ &\geq 2m_{i}-w_{i^{\prime}}>0\end{aligned}$
(recall that $x_{1}$ and $x_{2}$ are labels given by Definition 3.1 for the two elements of $S$, and $y$ is the single element of NN). Thus, if we consider $|A_{i}|$ and $f(A_{\leq i^{\prime}}\cup A_{i})$ as two potentials then no $(m_{i},\varepsilon)$-improvement ever decreases these potentials, and each such improvement either increases $|A_{i}|$ by 11 or increases $f(A_{\leq i^{\prime}}\cup A_{i})$ by at least $\varepsilon m_{i}$. Let us denote by $\hat{A}_{i}$ the final value of the set $A_{i}$ in iteration ii. Then, the potential $|A_{i}|$ is non-negative and upper bounded by $|\hat{A}_{i}|$, and therefore at most $|\hat{A}_{i}|$ $(m_{i},\varepsilon)$-improvements can increase it by 11. Similarly, by the submodularity of $f$, the potential $f(A_{\leq i^{\prime}}\cup A_{i})$ starts with a value of $f(A_{\leq i^{\prime}})$ and ends up with the value
$f(A_{\leq i^{\prime}}\cup\hat{A}_{i})\leq f(A_{\leq i^{\prime}})+\sum_{e\in\hat{A}_{i}}f(e\mid A_{\leq i^{\prime}})\leq f(A_{\leq i^{\prime}})+|\hat{A}_{i}|\cdot w_{i^{\prime}}\leq f(A_{\leq i^{\prime}})+2|\hat{A}_{i}|\cdot m_{i}\kern 5.0pt,$
where the last inequality employs Lemma A.2. Therefore, at most $2\varepsilon^{-1}|\hat{A}_{i}|$ improvements can increase this potential by $\varepsilon m_{i}$. Thus, the total number of $(m_{i},\varepsilon)$-improvements of both kinds in iteration $i$ of Algorithm 3 is upper bounded by $(1+2\varepsilon^{-1})|\hat{A}_{i}|$. Since the sets $A_{i}$ are kept disjoint by Algorithm 3, summing the last expression over all iterations of the algorithm yields $M\leq(1+2\varepsilon^{-1})|E|=O(\varepsilon^{-1}|E|)$. The proposition now follows by plugging this bound on $M$ into (6). ∎
Appendix B Analyzing the Distribution of $r_{o}$
The following lemma analyzes the distribution of $r_{o}$ for an arbitrary element $o\in O$. Recall that $O$ is an arbitrary strictly down-monotone feasible set.
Lemma B.1.
For each $o\in O$, the value $r_{o}$ defined in Section 4.5 is equal to $2^{\beta}$ for a random variable $\beta$ distributed uniformly at random in the range $[0,1)$.
Proof.
Recall that for all $i\geq 0$, $m_{i}=W\tau 2^{-i}=W2^{\alpha-i}$, where $\alpha$ is chosen uniformly at random from $(0,1]$, and note that $r_{o}=\min\{m_{i}/u(o)\mid m_{i}\geq u(o)\}$. Let
$i^{*}$ $=\lfloor\log_{2}W-\log_{2}u(o)\rfloor+1\kern 5.0pt,\quad\text{and}$
$\alpha^{*}$ $=i^{*}-(\log_{2}W-\log_{2}u(o))\kern 5.0pt.$
Then, we have $\alpha^{*}\in(0,1]$, and $i^{*}$ is a positive integer because the submodularity of $f$ guarantees that $u(o)\leq f(o\mid\emptyset)\leq W$. Additionally, since Observation 4.4 guarantees that $u(o)>0$, we get that $W$ must be positive as well. Therefore, it holds that $W2^{\alpha^{*}-i^{*}}=u(o)$, $W2^{-i^{*}}<u(o)$, and $W2^{-i^{*}+1}\geq u(o)$.
The above inequalities imply that, for all $\alpha\in(0,\alpha^{*})$,
$m_{i^{*}}$ $=W2^{\alpha-i^{*}}<W2^{\alpha^{*}-i^{*}}=u(o)\kern 5.0pt,\quad\text{and}$
$m_{i^{*}-1}$ $=W2^{\alpha-i^{*}+1}>W2^{-i^{*}+1}\geq u(o)\kern 5.0pt,$
and hence, $r_{o}=m_{i^{*}-1}/u(o)=(W2^{\alpha-i^{*}+1})/(W2^{\alpha^{*}-i^{*}})=2^{\alpha-\alpha^{*}+1}$. Similarly, for all $\alpha\in[\alpha^{*},1]$,
$m_{i^{*}}$ $=W2^{\alpha-i^{*}}\geq W2^{\alpha^{*}-i^{*}}=u(o)\kern 5.0pt,\quad\text{and}$
$m_{i^{*}+1}$ $=W2^{\alpha-i^{*}-1}\leq W2^{-i^{*}}<u(o)\kern 5.0pt,$
and hence, $r_{o}=m_{i^{*}}/u(o)=(W2^{\alpha-i^{*}})/(W2^{\alpha^{*}-i^{*}})=2^{\alpha-\alpha^{*}}$. Altogether, we then have $r_{o}=2^{\beta}$, where
$\beta=\begin{cases}\alpha-\alpha^{*}+1,&\alpha\in(0,\alpha^{*})\\ \alpha-\alpha^{*},&\alpha\in[\alpha^{*},1]\,.\end{cases}$
To complete the proof of the lemma, it only remains to show that $\beta$ is distributed uniformly at random in the range $[0,1)$. To show this, we note that for any value $\gamma\in[0,1-\alpha^{*}]$,
$\Pr[\beta\leq\gamma]=\Pr[\alpha^{*}\leq\alpha\leq\alpha^{*}+\gamma]=(\alpha^{*}+\gamma)-\alpha^{*}=\gamma\kern 5.0pt,$
and for any value $\gamma\in(1-\alpha^{*},1)$,
$\Pr[\beta\leq\gamma]={}$ $\Pr[\beta\leq 1-\alpha^{*}]+\Pr[1-\alpha^{*}<\beta\leq\gamma]$
$={}$ $1-\alpha^{*}+\Pr[0<\alpha\leq\gamma-1+\alpha^{*}]=1-\alpha^{*}+(\gamma-1+\alpha^{*})=\gamma\kern 5.0pt.\qed$